### v1

# Optimal e-value testing for properly constrained hypotheses

Abstract
Hypothesis testing via e-variables can be framed as a sequential betting game, where a player each round picks an e-variable. A good player’s strategy results in an effective statistical test that rejects the null hypothesis as soon as sufficient evidence arises. Building on recent advances, we address the question of restricting the pool of e-variables to simplify strategy design without compromising effectiveness. We extend the results of Clerico (2024), by characterising optimal sets of e-variables for a broad class of non-parametric hypothesis tests, defined by finitely many regular constraints. As an application, we discuss optimality in algorithmic mean estimation, including the case of heavy-tailed random variables.
1 Introduction
Hypothesis testing is the branch of statistics concerned with verifying if an observed data set is consistent with some given theory, typically referred to as the null hypothesis. Traditionally, this is formulated in terms of p-values, which represent the probability of obtaining an empirical test statistic at least as “extreme” as the one observed, under the assumption that the null hypothesis holds. However, most classical p-value methods are designed for single, fixed-sample experiments. In modern research, where data are often collected iteratively or experiments are repeated, recalculating p-values with additional data can lead to false positives and misleading conclusions. To overcome these limitations, recent developments propose using e-values as a more reliable alternative for sequential and adaptive hypothesis testing.
E-values quantify evidence against a hypothesis and allow flexible data collection strategies, such as optional stopping or continuation, and post hoc significance level tuning. While related ideas were developed decades ago by early works (Wald, 1945; Darling and Robbins, 1967; Lai, 1976; Siegmund, 1978), interest in e-values surged with a series of papers whose first preprints appeared over the space of few months in 2019 (Wasserman et al., 2020; Shafer, 2021; Vovk and Wang, 2021; Grünwald et al., 2024). Hypothesis testing with e-values can be phrased in terms of a sequential betting game (Ramdas et al., 2023), adopting a game-theoretic perspective on probability recently developed by Shafer and Vovk (2001, 2019). In this game, at each round the player selects an e-variable, namely a non-negative function with an expected value of at most one under the null hypothesis. The player then receives a reward, determined by the value of the picked e-variable at the next observed data point. If a significant grow of the cumulative reward is observed, the null hypothesis can be confidently rejected, as standard martingale analysis shows that this is unlikely to happen under the null hypothesis. We refer to Ramdas and Wang (2024) for an overview on e-value testing and game-theoretic interpretations.
A testing approach is effective if it leads to the rejection of a hypothesis as soon as there is enough evidence against it in the data. In our framework, this can be achieved if the player adopts a sound strategy, which yields high rewards whenever possible. The theoretical foundation for selecting such strategies in sequential games lies in online learning theory (Cesa-Bianchi and Lugosi, 2006; Orabona, 2023). To simplify strategy design and leverage online learning theoretical tools, it is often convenient to restrict the pool of e-variables among which the player can pick each round. In order to make this point more explicit, Clerico (2024) introduced the concept of restricted testing game. A key question is whether such restriction is detrimental for the test, by preventing the selection of strategies that could yield higher rewards. The answer is no, provided the restriction is well chosen. Moreover, in some cases, there exists an optimal class of e-variables to which one should restrict the game, informally the smallest set within which any good strategy must choose the e-variables. These ideas were developed in Clerico (2024), in the context of a simple e-variable testing problem that was motivated by mean estimation for a random variable bounded in $[0,1]$. This work extends these results to a wider class of hypothesis testing problems, by characterising the optimal set of e-variables for non-parametric hypotheses defined by a finite number of suitably regular constraints. This class is broad enough to capture meaningful problems, while allowing us to use standard analysis techniques, without relying on advanced tools from functional analysis and measure theory.
Notions of optimality for single e-variables, such as log-optimality (Koolen and Daniel, 2022; Larsson et al., 2024; Grünwald et al., 2024) and admissibility (Ramdas et al., 2022), have been previously extensively explored in the sequential hypothesis testing literature. However, as highlighted by Clerico (2024), the key distinction in our approach is the emphasis on optimality at the level of sets of e-variables, rather than focusing on individual e-variables.
1.1 Structure
This paper is organised as follows. In Section 2, we briefly review the fundamental standard concepts related to e-value hypothesis testing. Section 3 introduces the key definitions and basic properties of majorising and optimal classes of e-variables, while Section 4 defines the concept of properly constrained hypothesis, which serves as the central framework for this work. Section 5 discusses the dual class of e-variables for properly constrained hypotheses and states our main result: the optimality of the dual class (Theorem 1), which will be proved throughout the subsequent sections. Specifically, we begin by addressing the case of a finite sample space in Section 6, and then introduce the notion of matching sets in Section 7, which will be an essential tool to extend the optimality results to general closed sample spaces in Section 8. In Section 9, we broaden our results, by addressing tests of finitely (but not properly) constrained hypotheses, and by considering hypotheses defined via inequality constraints. In Section 10, our results are applied to the problem of deriving confidence sequences for the mean of (bounded and heavy-tailed) random variables. Lastly, Section 11 provides concluding remarks and outlines potential directions for future research.
1.2 Notation
Before delving into the main discussion, we first introduce the key notations and conventions used in this paper. We will often make use of topological notions (e.g., continuity, closedness, compactness, Borel measurability) on $\mathbb{R}^{n}$. The underlying topology is always implicitly assumed to be the standard Euclidean one. Throughout the whole paper, $\mathcal{X}$ denotes a closed set in $\mathbb{R}^{n}$.
For an arbitrary set $\mathcal{Y}\subseteq\mathbb{R}^{n}$, we let $\mathcal{P}_{\mathcal{Y}}$ be the set of Borel probability measures on $\mathbb{R}^{n}$ whose support is included in $\mathcal{Y}$. If $\mathcal{X}\subseteq\mathbb{R}^{n}$ is closed, with a slight abuse of notation we can identify $\mathcal{P}_{\mathcal{X}}$ with the set of all Borel probability measures on $\mathcal{X}$ (with respect to the induced topology). For $P\in\mathcal{P}_{\mathcal{Y}}$ and $\phi:\mathcal{Y}\to[0,+\infty)$ a Borel function, $\langle P,\phi(Y)\rangle$ (or more compactly $\langle P,\phi\rangle$) denotes the expectation of $\phi$ under $Y\sim P$. For $y\in\mathcal{Y}$, $\delta_{y}$ is the Dirac mass on $y$, namely $\langle\delta_{y},\phi\rangle=\phi(y)$. For $\Phi:\mathcal{Y}\to\mathbb{R}^{m}$, let $\|\Phi\|_{1}:\mathcal{Y}\to\mathbb{R}$ denote the mapping $y\mapsto\|\Phi(y)\|_{1}$, with $\|\cdot\|_{1}$ the 1-norm in $\mathbb{R}^{m}$. If $\Phi$ is Borel and $\langle P,\|\Phi\|_{1}\rangle$ is finite, then the expectation of $\Phi$ under $P$ is well defined (in $\mathbb{R}^{m}$) and we still denote it as $\langle P,\Phi\rangle$. As we will often define classes of probability measures in terms of the expectation with respect to some fixed Borel function $\Phi:\mathcal{Y}\to\mathbb{R}^{m}$, it is convenient to introduce the notation $\mathcal{P}_{\Phi}$ for the set of all measures in $\mathcal{P}_{\mathcal{Y}}$ for which $\langle P,\|\Phi\|_{1}\rangle$ is finite.
For $S\subseteq\mathcal{X}$, and $\Phi:\mathcal{X}\to\mathbb{R}^{m}$, we let $\Phi\big{|}_{S}:S\to\mathbb{R}^{m}$ be the restriction of $\Phi$ to $S$. We denote as $v\cdot u$ the scalar product when $u$ and $v$ are Euclidean vectors. Hence, for $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ and $\lambda\in\mathbb{R}^{m}$, $\lambda\cdot\Phi$ is a real function $\mathcal{X}\to\mathbb{R}$, mapping $x$ to $\lambda\cdot\Phi(x)$. We denote as $\operatorname*{Span}\Phi$ the linear span of the set $\{\Phi_{1},\dots,\Phi_{m}\}$ in the vector space of real functions on $\mathcal{X}$ (where $\Phi_{i}$ is the $i$-th component of $\Phi$), namely $\operatorname*{Span}\Phi=\{\lambda\cdot\Phi\,,\,\lambda\in\mathbb{R}^{m}\}$. For $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ and $S\subseteq\mathcal{X}$, $\Phi(S)$ is the image of $S$ under $\Phi$, namely $\{\Phi(x)\,:\,x\in S\}$. In particular, $\Phi(\mathcal{X})$ refers to the image of $\Phi$.
Given a set $\mathcal{Y}\subseteq\mathbb{R}^{n}$, we denote as $\operatorname*{conv}\mathcal{Y}$ its convex hull, and as $\operatorname*{aff}\mathcal{Y}$ its affine hull. We use the notation $\operatorname*{lin}\mathcal{Y}$ for the linear hull (span) of $\mathcal{Y}$. We denote as $\operatorname*{int}\mathcal{Y}$ the interior of $\mathcal{Y}$ and as $\operatorname*{rel\,int}\mathcal{Y}$ its relative interior. See Section B.2 for definitions from elementary convex analysis.
Sequences are denoted as $(x_{t})_{t\geq T_{0}}$, where $t$ is an integer-valued index and $T_{0}$ its smallest value. For high-probability statements, $\mathbb{P}$ expresses probability with respect to all the involved randomness. For instance, if $(X_{t})_{t\geq 1}$ is a sequence of independent draws from $P\in\mathcal{P}_{\mathcal{X}}$, we can write $\mathbb{P}\big{(}X_{t}\geq 0\,,\,\forall t\geq 1\big{)}$, with obvious meaning.
2 Algorithmic hypothesis testing
For the whole paper, $\mathcal{X}$ denotes a closed set in $\mathbb{R}^{n}$, endowed with the Borel sigma-field. We let $\mathcal{P}_{\mathcal{X}}$ be the set of Borel probability measures on $\mathcal{X}$.
Definition 1 (Hypotheses and e-variables).
A hypothesis (on $\mathcal{X}$) is a non-empty subset $\mathcal{H}$ of $\mathcal{P}_{\mathcal{X}}$. Given a hypothesis $\mathcal{H}$, an e-variable (with respect to $\mathcal{H}$) is a non-negative Borel function $E:\mathcal{X}\to[0,+\infty)$, such that $\langle P,E\rangle\leq 1$ for any $P\in\mathcal{H}$. We denote as $\mathcal{E}_{\mathcal{H}}$ the set of all the e-variables with respect to $\mathcal{H}$. We call e-class any subset of $\mathcal{E}_{\mathcal{H}}$.
Over the last few years, e-variables have become a fundamental tool in hypothesis testing, as they provide a practical and flexible alternative to the traditional p-value-based methods, allowing to overcome several known limitations of conventional testing procedures (Wang and Ramdas, 2022; Grünwald et al., 2024; Grünwald, 2024). E-variable testing is usually framed in terms of a sequential betting game, where at each round a player picks an e-variable (Ramdas and Wang, 2024). Using the terminology of Clerico (2024), we define the following testing game.
Definition 2 (Testing game).
Fix a hypothesis $\mathcal{H}$ on $\mathcal{X}$, and an e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. We call $\mathcal{E}$-restricted testing game (on $\mathcal{H}$) the following sequential procedure. Each round $t\geq 1$, a player
  • measurably [^2] We assume that the choice of $E_{t}$ is measurable, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable. picks an e-variable $E_{t}\in\mathcal{E}$, based solely on the past observations $x_{1},\dots,x_{t-1}$;
  • observes a new data point $x_{t}\in\mathcal{X}$;
  • earns a reward $\log E_{t}(x_{t})\in[-\infty,+\infty)$.
If $\mathcal{E}=\mathcal{E}_{\mathcal{H}}$, we speak of unrestricted testing game.
If the data points observed during the game are independent draws from $P\in\mathcal{H}$, the player’s cumulative reward is unlikely to grow excessively. This is formalised by the next statement, whose proof (see, e.g., Ramdas et al., 2023; Ramdas and Wang, 2024; Clerico, 2024) follows directly from Ville’s inequality (Ville, 1939), a uniform upper bound for non-negative super-martingales.
Lemma 1.
Let $\mathcal{H}$ be a hypothesis on $\mathcal{X}$ and consider a sequence $(X_{t})_{t\geq 1}\subseteq\mathcal{X}$ of independent draws from some $P\in\mathcal{H}$. Fix $\delta\in(0,1)$ and $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. Consider an $\mathcal{E}$-restricted testing game on $\mathcal{X}$, where the player observes the sequence $(X_{t})_{t\geq 1}$. Denote the player’s cumulative reward at round $T$ as $R_{T}=\sum_{t=1}^{T}\log E_{t}(X_{t})$. Then, we have
$\mathbb{P}\big{(}R_{T}\leq\log\tfrac{1}{\delta}\,,\;\forall T\geq 1\big{)}\geq 1-\delta\,.$
The cumulative reward of the player can be interpreted as evidence accumulated against $\mathcal{H}$. Specifically, Lemma 1 justifies a sequential testing procedure, where the null hypothesis (asserting that the data are generated from a $P\in\mathcal{H}$) is rejected whenever the player’s cumulative reward exceeds the threshold value of $\log(1/\delta)$. Here, $\delta\in(0,1)$ denotes the type II confidence level, meaning that if the null hypothesis is true, it will be rejected with probability at most $\delta$. It is worth noticing that the high-probability inequality in Lemma 1 holds uniformly for any data set size $T$. This ensures that when data are collected sequentially, optional stopping or continuation are allowed, an advantage not afforded by standard p-value methods. For a comprehensive overview of sequential testing by betting, we refer to Chapter 6 of Ramdas and Wang (2024).
It is clear that a player always selecting the constant function $1$ will never reject a hypothesis, as its cumulative reward is always null. To get an effective test, we want the cumulative reward to be “as large as possible”, so that the null hypothesis is rejected whenever there is sufficient evidence against it in the observed data. It is crucial to carefully determine the pool $\mathcal{E}$ of e-variables from which the player can pick: a poor choice may exclude useful e-variables, while there is no advantage in keeping redundant ones (e.g., the constant function $1/2$). A well chosen $\mathcal{E}$ can simplify the problem and enhance the design of effective betting strategy, without compromising the strength of the test. This point was addressed by Clerico (2024), with the introduction of the notions of majorising and optimal e-classes, which are the subject of the next section.
3 Majorising e-classes and optimal e-class
We now present the definitions and basic properties of majorising and optimal e-classes, as introduced in Clerico (2024). In the following, fix a closed $\mathcal{X}\subseteq\mathbb{R}^{n}$ and a hypothesis $\mathcal{H}$ on $\mathcal{X}$. We consider the standard partial ordering on the set of real functions on $\mathcal{X}$. Specifically, given two functions $f$ and $f^{\prime}$ from $\mathcal{X}$ to $\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in\mathcal{X}$. If $f\succeq f^{\prime}$ and $f\neq f^{\prime}$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$. We will also use the symbols $\preceq$ and $\prec$, with obvious meaning.
Definition 3 (Majorising and optimal e-classes).
An e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$ is a majorising e-class when, for any $E\in\mathcal{E}_{\mathcal{H}}$, we can find an e-variable $E^{\prime}\in\mathcal{E}$ such that $E^{\prime}\succeq E$. If a majorising e-class is contained in every other majorising e-class, it is called optimal.
It is clear that $\mathcal{E}_{\mathcal{H}}$ is itself a majorising e-class. Conversely, an optimal e-class may not exist. If it does exist, it is unique and corresponds to the intersection of all majorising e-classes. Next, we provide a necessary and sufficient condition for the existence of the optimal e-class, based on the notion of maximality for e-variables. For the proof, see Lemmas 3 and 4 in Clerico (2024).
Definition 4 (Maximal e-variables).
A maximal e-variable is an e-variable $E\in\mathcal{E}_{\mathcal{H}}$ that has no strict majoriser, namely there is no $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succ E$.
Lemma 2.
Every majorising e-class includes all maximal e-variables. An optimal e-class exists if, and only if, there is a majorising e-class whose elements are all maximal. In particular, if an optimal e-class exists, it is unique and it corresponds to the set of all maximal e-variables.
Corollary 1.
If, for some $x_{0}\in\mathcal{X}$, $\sup_{P\in\mathcal{H}}P(\{x_{0}\})=0$, there is no optimal e-class for $\mathcal{H}$.
Proof.
Fix $E\in\mathcal{E}_{\mathcal{H}}$ and let $\mathbf{1}_{x_{0}}:\mathcal{X}\to\mathbb{R}$ be $1$ on $x_{0}$ and 00 everywhere else. Then, for any $P\in\mathcal{H}$, we have $\langle P,E+\mathbf{1}_{x_{0}}\rangle=\langle P,E\rangle\leq 1$, and so $E+\mathbf{1}_{x_{0}}\in\mathcal{E}_{\mathcal{H}}$. Since $E+\mathbf{1}_{x_{0}}\succ E$, $E$ is not maximal. We deduce that the set of all maximal e-variables is empty, and we conclude by Lemma 2. ∎
Majorising and optimal e-classes play a fundamental role in our analysis of sequential hypothesis testing. Restricting the game of Definition 2 to a majorising e-class does not hinder the player’s performance. To illustrate this, consider two players, Alice and Bob, observing the same sequence of data points. Bob plays an unrestricted testing game, while Alice’s choices are restricted to a majorising e-class $\mathcal{E}$. Assume that Bob makes his move first, selecting an e-variable $E_{t}$. Alice, aware of Bob’s choice, can then pick $E_{t}^{\prime}\in\mathcal{E}$ such that, for any possible observation $x_{t}$, $E_{t}^{\prime}(x_{t})\geq E_{t}(x_{t})$. Thus, Alice can opt for a strategy ensuring that her cumulative reward is always at least as large as Bob’s, even though her choices are restricted to $\mathcal{E}$. Now, suppose that $\mathcal{E}$ is not just majorising, but optimal. In this case, if Bob chooses an e-variable $E_{t}$ outside of $\mathcal{E}$, Alice can select an alternative $E_{t}^{\prime}\in\mathcal{E}$ such that $E_{t}^{\prime}\succ E_{t}$. Such $E_{t}^{\prime}$ provides rewards at least as good as those of $E_{t}$ for all possible values of the next observation, and strictly better for at least one value of $x_{t}$. On the other hand, if Alice picks first, the only way for Bob to be absolutely sure that his reward will not be worse than Alice’s is to select the same e-variable, and thus choose from the optimal e-class. This is because Alice’s choice is maximal (Lemma 2), meaning no other e-variable can majorise it. In this sense, the optimal e-class simplifies designing effective betting strategies, by eliminating all redundant and possibly detrimental e-variables.
We conclude this section with some positive result on the existence of the optimal e-class.
Lemma 3.
If $\mathcal{X}$ is finite and $\sup_{P\in\mathcal{H}}P(\{x\})>0$ for every $x\in\mathcal{X}$, the optimal e-class exists.
Proof.
Let $d<+\infty$ be the cardinality of $\mathcal{X}$, and write $\mathcal{X}=\{x_{1},\dots,x_{d}\}$. Any $E\in\mathcal{E}_{\mathcal{H}}$ and $P\in\mathcal{H}$ can be seen as elements of $\mathbb{R}^{d}$, with $\langle P,E\rangle$ being the standard dot product. For the rest of this proof we will always consider the standard Euclidean topology on $\mathbb{R}^{d}$ to define continuity, closedness and compactness. First, note that $\mathcal{E}_{\mathcal{H}}=\bigcap_{P\in\mathcal{H}}\{E\succeq 0\,:\,\langle P,E\rangle\leq 1\}$ is closed, as the intersection of closed sets. Moreover, it is also bounded, and thus compact. Indeed, for any $x\in\mathcal{X}$ and $E\in\mathcal{E}_{\mathcal{H}}$, we have $E(x)\sup_{P\in\mathcal{H}}P(\{x\})\leq\sup_{P\in\mathcal{H}}\langle P,E\rangle\leq 1$ and $\sup_{P\in\mathcal{H}}P(\{x\})>0$.
By Lemma 2, it is enough to show that the set of all maximal e-variables is a majorising e-class, or equivalently that every e-variable is majorised by a maximal e-variable. So, fix $E_{0}\in\mathcal{E}_{\mathcal{H}}$. The set $\mathcal{E}_{0}=\{E\in\mathcal{E}_{\mathcal{H}}\,:\,E\succeq E_{0}\}\subseteq\mathcal{E}_{\mathcal{H}}$ is closed, hence compact. Moreover, the mapping $E\mapsto E(x_{1})$ is continuous on $\mathcal{E}_{\mathcal{H}}$. We can thus find $E_{1}\in\operatorname*{arg\,max}_{E\in\mathcal{E}_{0}}E(x_{1})$. We iterate this procedure up to $i=d$, with $\mathcal{E}_{i}=\{E\in\mathcal{E}_{\mathcal{H}}\,:\,E\succeq E_{i}\}$ and $E_{i}\in\operatorname*{arg\,max}_{E\in\mathcal{E}_{i-1}}E(x_{i})$. Now, by construction $E_{d}\succeq E_{0}$. Consider any e-variable $E^{\star}\succeq E_{d}$ and fix any integer index $i\in[1,d]$. Then, $E^{\star}\succeq E_{d}\succeq E_{i-1}$, which by the definition of $E_{i}$ implies that $E^{\star}(x_{i})\leq E_{i}(x_{i})\leq E_{d}(x_{i})$, and so $E^{\star}(x_{i})=E_{d}(x_{i})$. Therefore, $E^{\star}=E_{d}$ and $E_{d}$ is maximal. ∎
Although we will not make use of it, we state next an extension of the lemma above, for countable $\mathcal{X}$. A proof, via the construction of a transfinite sequence, is detailed in Appendix A.
Lemma 4.
If $\mathcal{X}$ is countable and $\sup_{P\in\mathcal{H}}P(\{x\})>0$ for all $x\in\mathcal{X}$, the optimal e-class exists.
One might think that if $\sup_{P\in\mathcal{H}}P(\{x\})>0$ for every $x\in\mathcal{X}$ the optimal e-class always exists. However, this is not always the case and a counter example is provided in Appendix C.
4 Properly constrained hypotheses
In the remainder of this paper, we will mainly focus on a specific kind of hypotheses, for which we aim to fully characterise the optimal e-class. Notably, Clerico (2024) examined the simple case where $\mathcal{X}=[0,1]$ and $\mathcal{H}$ is the set of probability measures on $\mathcal{X}$ with a fixed mean. Here, we discuss the more general setting where $\mathcal{X}$ is a generic closed set in $\mathbb{R}^{n}$ and the hypothesis $\mathcal{H}$ is defined in terms of the expectation of a continuous function $\Phi:\mathcal{X}\to\mathbb{R}^{m}$. Since the approach in Clerico (2024) is specifically designed for their simpler setting, we will have to develop more sophisticated proof techniques to handle the broader framework that we consider in this work.
In the following, fix a closed set $\mathcal{X}\subseteq\mathbb{R}^{n}$. We recall that, for $\Phi:\mathcal{X}\to\mathbb{R}^{m}$, $\mathcal{P}_{\Phi}$ is the set of measures $P$ in $\mathcal{P}_{\mathcal{X}}$ such that $\langle P,\|\Phi\|_{1}\rangle<+\infty$.
Definition 5 (Constraints).
Let $\mathcal{H}$ be a hypothesis on $\mathcal{X}$. If a continuous $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ satisfies
=$\mathcal{H}=\{P\in\mathcal{P}_{\Phi}\;:\;\langle P,\Phi\rangle=0\}\,,$
we call $\Phi$ a constraint for $\mathcal{H}$. Let $\mathcal{C}$ denote the convex hull of $\Phi(\mathcal{X})$. If $0\in\operatorname*{rel\,int}\mathcal{C}$ we say that $\Phi$ is a proper constraint. If moreover $0\in\operatorname*{int}\mathcal{C}$, then we call $\Phi$ a minimal constraint.
Definition 6 (Finitely and properly constrained hypotheses).
A hypothesis is called finitely constrained if it admits a constraint. It is properly constrained if it admits a proper constraint.
We remark that any finitely constrained $\mathcal{H}$ is convex. Indeed, $\mathcal{P}_{\Phi}$ is convex, and $\mathcal{H}$ is a level set of the convex functional $P\mapsto\langle P,\Phi\rangle$ on $\mathcal{P}_{\Phi}$. Next, we list a few basic properties of properly constrained hypotheses. The proofs, detailed in Appendix A, are generally rather short and rely on elementary tools from linear algebra and convex analysis.
Lemma 5.
Every properly constrained hypothesis admits a minimal constraint.
Lemma 6.
Let $\mathcal{H}$ be a properly constrained hypothesis and $\Phi$ a proper constraint. $\Phi$ is minimal if, and only if, all its components are linearly independent scalar functions.
Lemma 7.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$. Then, for every $x\in\mathcal{X}$ there exists a $P\in\mathcal{H}$ whose support has finite cardinality and such that $x\in\operatorname*{Supp}P$.
Lemma 8.
Let $\mathcal{H}$ be a finitely constrained hypothesis on $\mathcal{X}$. $\mathcal{H}$ is properly constrained if, and only if, $\sup_{P\in\mathcal{H}}P(\{x\})>0$ for all $x\in\mathcal{X}$. Moreover, if $\mathcal{H}$ is properly constrained, every constraint $\Phi$ for $\mathcal{H}$ is a proper constraint and there is $A:\mathcal{X}\to[1,+\infty)$ such that $E\preceq A$ for any $E\in\mathcal{E}_{\mathcal{H}}$.
Lemma 9.
Let $\mathcal{H}$ be a finitely constrained hypothesis on $\mathcal{X}$. $\mathcal{H}$ is properly constrained if, and only if, there is $P\in\mathcal{H}$ such that $\operatorname*{Supp}P=\mathcal{X}$.
As a side note, we state a stronger version of the last claim in Lemma 8, which may be of future interest, while not directly used in this work. The proof is deferred to Appendix D.
Proposition 1.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$. There is a continuous function $A:\mathcal{X}\to[1,+\infty)$ such that $\langle P,A\rangle<+\infty$ for every $P\in\mathcal{H}$, $\sup_{P\in\mathcal{H}}P(\{x\})\geq 1/A(x)$ for all $x\in\mathcal{X}$, and $E\preceq A$ for any $E\in\mathcal{E}_{\mathcal{H}}$.
5 Dual e-class
For any properly constrained hypothesis $\mathcal{H}$ on a closed $\mathcal{X}\subseteq\mathbb{R}^{n}$, we can define a dual e-class, which will play a fundamental role in our analysis. Indeed, it will turn out that it coincides precisely with the optimal e-class. First, we introduce the following notation, which will be used throughout the rest of the paper. For a function $\Phi:\mathcal{X}\to\mathbb{R}^{m}$, we let
$\Lambda_{\Phi}=\left\{\lambda\in\mathbb{R}^{m}\;:\;\sup_{x\in\mathcal{X}}\lambda\cdot\Phi(x)\leq 1\right\}\,.$
Definition 7.
Let $\mathcal{H}$ be a properly constrained hypothesis and $\Phi$ a constraint. The dual e-class of $\mathcal{H}$ is
=$\mathcal{E}_{\mathcal{H}}^{\vee}=\left\{1-\lambda\cdot\Phi\;:\;\lambda\in\Lambda_{\Phi}\right\}\,.$
It is easy to check that $\mathcal{E}_{\mathcal{H}}^{\vee}$ is an e-class. Indeed, any $E\in\mathcal{E}_{\mathcal{H}}^{\vee}$ is non-negative (by the definition of $\Lambda_{\Phi}$), continuous, and $\langle P,E\rangle=1$ for all $P\in\mathcal{H}$. Moreover, $\mathcal{E}_{\mathcal{H}}^{\vee}$ is well defined, in the sense that it is independent of the specific constraint $\Phi$ used in its definition. To see this, recall that the span of $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ is $\operatorname*{Span}\Phi=\{\lambda\cdot\Phi\,:\,\lambda\in\mathbb{R}^{m}\}$, and note that =$\mathcal{E}_{\mathcal{H}}^{\vee}=\{1-\phi\succeq 0\,:\,\phi\in\operatorname*{Span}\Phi\}$. The next result implies that this set is the same for any constraint $\Phi$ for $\mathcal{H}$.
Proposition 2.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$, with $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ and $\Phi^{\prime}:\mathcal{X}\to\mathbb{R}^{m^{\prime}}$ two constraints. Then, $\operatorname*{Span}\Phi=\operatorname*{Span}\Phi^{\prime}$.
When $\mathcal{X}$ has finite cardinality, Proposition 2 is a direct consequence of Lemma 11 (stated in the next section). For a general closed $\mathcal{X}\subseteq\mathbb{R}^{n}$, we defer the proof to the end of Section 7.
We now state the main result of this work: for any properly constrained hypothesis, the optimal e-class exists and coincides with the dual e-class.
Theorem 1.
Let $\mathcal{H}$ be a properly constrained hypothesis. The dual e-class is the optimal e-class.
We will first prove the optimality of the dual e-class for the case where $\mathcal{X}$ has finite cardinality (Proposition 3). The notion of matching set, which we will introduce in Section 7, will enable us to build on this preliminary result to establish the optimality of $\mathcal{E}_{\mathcal{H}}^{\vee}$ for compact $\mathcal{X}$ (Proposition 4). Finally, we will prove the general case of a closed $\mathcal{X}$ at the end of Section 8.
For now, we start by checking that all the elements of $\mathcal{E}_{\mathcal{H}}^{\vee}$ are maximal.
Lemma 10.
Let $\mathcal{H}$ be a properly constrained hypothesis. Every $E\in\mathcal{E}_{\mathcal{H}}^{\vee}$ is maximal.
Proof.
Fix $E\in\mathcal{E}_{\mathcal{H}}^{\vee}$, and consider an e-variable $\hat{E}\succeq E$. Fix any $x\in\mathcal{X}$. By Lemma 8, there is a $P\in\mathcal{H}$ with an atom on $x$, and the fact that $E\in\mathcal{E}_{\mathcal{H}}^{\vee}$ implies that $\langle P,E\rangle=1$. Hence, $P(\{x\})(\hat{E}(x)-E(x))\leq\langle P,\hat{E}-E\rangle=\langle P,\hat{E}\rangle-1\leq 0$, because $\hat{E}$ is an e-variable and $\hat{E}-E$ is non-negative. Since $P(\{x\})>0$ and $\hat{E}\succeq E$, it must be that $\hat{E}(x)=E(x)$. So, $E$ is maximal. ∎
6 Domains with finite cardinality
For this section, we let $\mathcal{X}$ have finite cardinality $d$. Therefore, we can see probability measures $P$ on $\mathcal{X}$ and scalar functions $\phi$ on $\mathcal{X}$ as vectors in $\mathbb{R}^{d}$, so that $\langle P,\phi\rangle$ is now simply the dot product in $\mathbb{R}^{d}$. With this in mind, a hypothesis $\mathcal{H}$ on $\mathcal{X}$ is a subset of $\mathbb{R}^{d}$, and we can define $\mathcal{H}^{\perp}$ as the usual orthogonal set in $\mathbb{R}^{d}$. Moreover, we remark that, for any $\Phi:\mathcal{X}\to\mathbb{R}^{m}$, $\mathcal{P}_{\Phi}=\mathcal{P}_{\mathcal{X}}$.
Lemma 11.
Let $\mathcal{X}$ have finite cardinality. Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$ and $\Phi$ a constraint. Then, $\operatorname*{Span}\Phi=\mathcal{H}^{\perp}$.
Proof.
First, note that $\mathcal{H}=(\operatorname*{Span}\Phi)^{\perp}\cap\mathcal{P}_{\mathcal{X}}$. By Lemma 9 and Lemma 25, $\mathcal{H}\cap\operatorname*{rel\,int}\mathcal{P}_{\mathcal{X}}\neq\varnothing$. In particular, by Lemma 26, $\operatorname*{lin}\mathcal{H}=\operatorname*{lin}((\operatorname*{Span}\Phi)^{\perp}\cap\mathcal{P}_{\mathcal{X}})=(\operatorname*{Span}\Phi)^{\perp}$. As we are dealing with finite dimensional spaces, $\operatorname*{Span}\Phi=(\operatorname*{Span}\Phi)^{\perp\perp}=(\operatorname*{lin}\mathcal{H})^{\perp}=\mathcal{H}^{\perp}$. ∎
Proposition 3.
If $\mathcal{H}$ is a properly constrained hypothesis on a finite set $\mathcal{X}$, $\mathcal{E}_{\mathcal{H}}^{\vee}$ is optimal.
Proof.
Lemma 8 and Lemma 3 ensure that the optimal e-class exists. Hence, by Lemma 2, it is sufficient to show that $\mathcal{E}_{\mathcal{H}}^{\vee}$ coincides with the set of all maximal e-variables. By Lemma 10, we know that all the elements of $\mathcal{E}_{\mathcal{H}}^{\vee}$ are maximal. Therefore, we are left to show that every maximal e-variable is in $\mathcal{E}_{\mathcal{H}}^{\vee}$. Fix a maximal e-variable $E$. We will adopt the following proof strategy. First, we show that there is a fully supported $\tilde{P}\in\mathcal{H}$ such that $\langle\tilde{P},E\rangle=1$. Using that such $\tilde{P}$ is in the relative interior of $\mathcal{H}$, we will establish that $\langle P,E\rangle=1$ for all $P\in\mathcal{H}$. Finally, we will see that this property is enough to ensure that $E\in\mathcal{E}_{\mathcal{H}}^{\vee}$.
We start by noticing that since $\mathcal{X}$ has finite cardinality, $\mathcal{H}$ is a convex and closed polytope (it is the intersection of a finite dimensional closed simplex with an affine space). Letting $\mathcal{V}\subseteq\mathcal{H}$ denote the (finite) set of vertices of $\mathcal{H}$, every $P\in\mathcal{H}$ is a convex combination of elements of $\mathcal{V}$.
Fix a maximal e-variable $E$, [^3] Note that a maximal e-variable has to exist by Lemma 2, since we already know that the optimal e-class exists. A more direct argument: the constant function $1$ belongs to $\mathcal{E}_{\mathcal{H}}^{\vee}$, by Lemma 10 it is a maximal e-variable. and let $\mathcal{V}_{0}=\{\hat{P}\in\mathcal{V}\,:\,\langle\hat{P},E\rangle=1\}$. $\mathcal{V}_{0}$ is non-empty. Indeed, $E$ cannot be identically null, so Lemma 8 implies that $C=\sup_{P\in\mathcal{H}}\langle P,E\rangle>0$. Hence, $E/C$ is still an e-variable. Since $C\leq 1$, the maximality of $E$ yields $C=1$. Therefore, $1=\sup_{P\in\mathcal{H}}\langle P,E\rangle=\max_{\hat{P}\in\mathcal{V}}\langle\hat{P},E\rangle$, since every $P\in\mathcal{H}$ is a convex combination of vertices of $\mathcal{H}$. Hence, $\mathcal{V}_{0}\neq\varnothing$.
Let $S=\bigcup_{\hat{P}\in\mathcal{V}_{0}}\operatorname*{Supp}\hat{P}$. Now, we will show that $S=\mathcal{X}$. By contradiction, let us assume that this is not the case, and fix $x\in\mathcal{X}\setminus S$. Let $\mathbf{1}_{x}$ denote the function that is $1$ on $x$ and 00 everywhere else. If $S\neq\mathcal{X}$, then $\mathcal{V}_{0}\neq\mathcal{V}$. So, we can define the mapping $\varphi_{x}:[0,+\infty)\to\mathbb{R}$ as
$\varphi_{x}\;:\;\varepsilon\mapsto\max_{\hat{P}\in\mathcal{V}\setminus\mathcal{V}_{0}}\langle\hat{P},E+\varepsilon\mathbf{1}_{x}\rangle\,.$
Since this is the maximum among a finite number of continuous functions, it is continuous. For $\varepsilon$ large enough, $\varphi_{x}(\varepsilon)>1$. Since $\varphi_{x}(0)<1$, there is $\tilde{\varepsilon}>0$ such that $\varphi_{x}(\tilde{\varepsilon})=1$. Let $\tilde{E}=E+\tilde{\varepsilon}\mathbf{1}_{x}$. Clearly, for any $\hat{P}\in\mathcal{V}_{0}$ we have $\langle\hat{P},\tilde{E}\rangle=\langle\hat{P},E\rangle=1$, since $x\notin\operatorname*{Supp}(\hat{P})$. Moreover, for all $\hat{P}\in\mathcal{V}\setminus\mathcal{V}_{0}$, $\langle\hat{P},\tilde{E}\rangle\leq 1$, by definition of $\varphi_{x}$ and $\tilde{\varepsilon}$. Therefore, $\tilde{E}$ is an e-variable, since its expectation is at most one under all the vertices of $\mathcal{H}$. As $\tilde{\varepsilon}>0$, $\tilde{E}\succ E$, and so $E$ cannot be maximal, which is a contradiction. We conclude that $S=\mathcal{X}$.
From what we have shown so far, there is a fully supported $P^{\star}\in\mathcal{H}$ such that $\langle P^{\star},E\rangle=1$. Indeed, it is enough to take $P^{\star}=N^{-1}\sum_{\hat{P}\in\mathcal{V}_{0}}\hat{P}$, where $N$ is the cardinality of $\mathcal{V}_{0}$. In particular, by Lemma 25, $P^{\star}\in\operatorname*{rel\,int}\mathcal{P}_{\mathcal{X}}$. Since $\mathcal{H}=\mathcal{P}_{\mathcal{X}}\cap(\operatorname*{Span}\Phi)^{\perp}$, and $\operatorname*{rel\,int}(\mathcal{P}_{\mathcal{X}})\cap(\operatorname*{Span}\Phi)^{\perp}$ is non-empty ($P^{\star}$ belongs to it), $\operatorname*{rel\,int}\mathcal{H}=\operatorname*{rel\,int}(\mathcal{P}_{\mathcal{X}})\cap(\operatorname*{Span}\Phi)^{\perp}$ (see, e.g., Proposition 2.1.10 in Chapter A of Hiriart-Urrut and Lemaréchal, 2004). So, $P^{\star}\in\operatorname*{rel\,int}\mathcal{H}$. As $\mathcal{H}$ is a convex polytope, there are strictly positive coefficients $(\alpha_{\hat{P}})_{\hat{P}\in\mathcal{V}}$, summing to $1$, such that $P^{\star}=\sum_{\hat{P}\in\mathcal{V}}\alpha_{\hat{P}}\hat{P}$ (see, e.g., Remark 2.1.4 in Chapter A of Hiriart-Urrut and Lemaréchal, 2004). In particular,
$0=\langle P^{\star},E\rangle-1=\sum_{\hat{P}\in\mathcal{V}}\alpha_{\hat{P}}(\langle\hat{P},E\rangle-1)\,,$
where $\alpha_{\hat{P}}>0$ for all $\hat{P}\in\mathcal{V}$, and so $\langle\hat{P},E\rangle=1$ for all $\hat{P}\in\mathcal{V}$. Hence, $\langle P,E\rangle=1$ for all $P\in\mathcal{H}$.
Now, let $\Phi$ be a constraint for $\mathcal{H}$. $\langle P,E\rangle=1$ for all $P\in\mathcal{H}$, so $E-1\in\mathcal{H}^{\perp}=\operatorname*{Span}\Phi$ by Lemma 11. Since $E$ is non-negative, $E=1+\lambda\cdot\Phi$ for some $\lambda\in\Lambda_{\Phi}$, and so $E\in\mathcal{E}_{\mathcal{H}}^{\vee}$. ∎
7 Matching sets
Thus far, we have shown the optimality of the dual e-class when $\mathcal{X}$ has finite cardinality. The notions of compatible and matching sets that we now introduce will allow us to build on these finite-cardinality results to address the case of a generic closed set $\mathcal{X}\subseteq\mathbb{R}^{n}$.
For any $S\subseteq\mathcal{X}$, recall that $\mathcal{P}_{S}=\{P\in\mathcal{P}_{\mathcal{X}}\,:\,\operatorname*{Supp}P\subseteq S\}$. For a measurable $\Phi:\mathcal{X}\to\mathbb{R}^{m}$, let $\mathcal{P}_{\Phi|S}=\{P\in\mathcal{P}_{S}\,:\,\langle P,\|\Phi\|_{1}\rangle<+\infty\}$. We notice that $\mathcal{P}_{\Phi|S}=\mathcal{P}_{\Phi}\cap\mathcal{P}_{S}$.
Definition 8 (Compatible and matching sets).
Let $\mathcal{H}$ be a hypothesis on $\mathcal{X}$ and $S\subseteq\mathcal{X}$ a closed set. We say that $S$ is a compatible set (for $\mathcal{H}$) if there is $P\in\mathcal{H}$ such that $\operatorname*{Supp}P\subseteq S$. We say that $S$ is a matching set (for $\mathcal{H}$) if there is $P\in\mathcal{H}$ such that $\operatorname*{Supp}P=S$.
Definition 9 (Restriction of a hypothesis).
The restriction of a hypothesis $\mathcal{H}$ on $\mathcal{X}$ to a compatible set $S\subseteq\mathcal{X}$ is defined as
$\mathcal{H}_{S}=\mathcal{H}\cap\mathcal{P}_{S}=\{P\in\mathcal{H}\;:\;\operatorname*{Supp}P\subseteq S\}\,.$
By definition of compatible set, $\mathcal{H}_{S}$ is non-empty, and hence a hypothesis on $S$. We list here a few results outlining basic properties of the restrictions of finitely and properly constrained hypotheses. The proofs, rather short, can be found in Appendix A.
Lemma 12.
Let $\mathcal{H}$ be a finitely constrained hypothesis on $\mathcal{X}$ and $\Phi$ a constraint. A closed set $S\subseteq\mathcal{X}$ is compatible if, and only if, $0\in\operatorname*{conv}\Phi(S)$. In such case, $\mathcal{H}_{S}$ is a finitely constrained hypothesis on $S$ and $\Phi\big{|}_{S}$ is a constraint for it. In particular, $\mathcal{H}_{S}=\{P\in\mathcal{P}_{\Phi|S}\,:\,\langle P,\Phi\big{|}_{S}\rangle=0\}$.
Lemma 13.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$ and $S\subseteq\mathcal{X}$ a compatible set. $S$ is a matching set if, and only if, $\mathcal{H}_{S}$ is a properly constrained hypothesis on $S$.
Lemma 14.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$. A finite union of matching sets is a matching set. In particular, every finite subset of $\mathcal{X}$ is contained in a finite matching set.
As a first application of what we have just introduced, we can now prove Proposition 2.
Proposition 2.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$, with $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ and $\Phi^{\prime}:\mathcal{X}\to\mathbb{R}^{m^{\prime}}$ two constraints. Then, $\operatorname*{Span}\Phi=\operatorname*{Span}\Phi^{\prime}$.
Proof.
If $\mathcal{X}$ is finite, the result is a direct consequence of Lemma 11. So, consider the case where $\mathcal{X}$ has infinitely many elements. Clearly, it is enough to show that $\operatorname*{Span}\Phi^{\prime}\subseteq\operatorname*{Span}\Phi$, or equivalently that, for any $\phi\in\operatorname*{Span}\Phi^{\prime}$, $\operatorname*{Span}\Phi\cup\{\phi\}$ is not a linearly independent family.
Fix $\phi\in\operatorname*{Span}\Phi^{\prime}$. Let $d$ be the dimension of $\operatorname*{Span}\Phi$. By Lemma 20, it is enough to show that for any $S\subseteq\mathcal{X}$ with $d+1$ elements, $V_{S}=\operatorname*{lin}(\operatorname*{Span}\Phi\big{|}_{S}\cup\{\phi\big{|}_{S}\})$ has dimension at most $d$. So, fix any $S\subseteq\mathcal{X}$ with $d+1$ elements. By Lemma 14, there is a finite matching set $S^{\prime}\supseteq S$. Since $\mathcal{H}_{S^{\prime}}$ is a properly constrained hypothesis on $S^{\prime}$ (Lemma 13) and both $\Phi\big{|}_{S^{\prime}}$ and $\Phi^{\prime}\big{|}_{S^{\prime}}$ are constraints for it (Lemma 12), we have that $\operatorname*{Span}\Phi\big{|}_{S^{\prime}}=\operatorname*{Span}\Phi^{\prime}\big{|}_{S^{\prime}}$ by Lemma 11, because $S^{\prime}$ has finite cardinality. As $S\subseteq S^{\prime}$, $\phi\big{|}_{S}\in\operatorname*{Span}\Phi\big{|}_{S}$. So, $\dim V_{S}=\dim(\operatorname*{Span}\Phi\big{|}_{S})\leq\dim(\operatorname*{Span}\Phi)=d$. ∎
8 Optimality of the dual e-class
At the end of this section we finally prove Theorem 1. First, we will establish the result for the case where $\mathcal{X}$ is compact, then extend it to the general case of a closed $\mathcal{X}\subseteq\mathbb{R}^{n}$. We introduce the following notation. For any $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ and any $S\subseteq\mathcal{X}$, we define the set
$\Lambda_{\Phi|S}=\left\{\lambda\in\mathbb{R}^{m}\;:\;\sup_{x\in S}\lambda\cdot\Phi(x)\leq 1\right\}\,.$
We state here two elementary properties of these sets.
Lemma 15.
For any sequence $(S_{i})_{i\geq 1}$ of subsets of $\mathcal{X}$, $\bigcap_{i=1}^{\infty}\Lambda_{\Phi|S_{i}}=\Lambda_{\Phi|S}$, with $S=\bigcup_{i=1}^{\infty}S_{i}$.
Proof.
For all $i\geq 1$, $S\supseteq S_{i}$, so $\Lambda_{\Phi|S}\subseteq\bigcap_{i=1}^{\infty}\Lambda_{\Phi|S_{i}}$. Moreover, if $\lambda\in\bigcap_{i=1}^{\infty}\Lambda_{\Phi|S_{i}}$, then we have $\lambda\cdot\Phi(x)\leq 1$ for every $x\in S_{i}$, for any $i\geq 1$. Therefore, $\lambda\cdot\Phi(x)\leq 1$ for all $x\in S$, so $\lambda\in\Lambda_{\Phi|S}$. ∎
Lemma 16.
If $\Phi$ is continue, $\Lambda_{\Phi|S}$ is closed, for any $S\subseteq\mathcal{X}$.
Proof.
It follows from the lower semi-continuity of the supremum of continuous functions. ∎
From Lemma 16, if $\Phi$ is a constraint of a finitely generated hypothesis, $\Lambda_{\Phi}$ is closed. If $\Phi$ is a minimal constraint of a properly generated hypothesis, a stronger result holds: $\Lambda_{\Phi}$ is compact.
Lemma 17.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$ and $\Phi$ a constraint. Then $\Phi$ is minimal, if, and only if, $\Lambda_{\Phi}$ is compact. More precisely, $\Phi$ is minimal if, and only if, there is a finite matching set $S^{\star}$ such that $\Lambda_{\Phi|S^{\star}}$ is compact. If such $S^{\star}$ exists, $\Lambda_{\Phi|S}$ is compact for any $S\subseteq\mathcal{X}$ that contains $S^{\star}$, and any closed set $S\subseteq\mathcal{X}$ that contains $S^{\star}$ is a matching set.
The proof of Lemma 17 (detailed in Appendix A) builds on the characterisation of minimal constraints that we stated in Lemma 6. An immediate corollary of the above is the following.
Corollary 2.
If $\mathcal{H}$ is a properly constrained hypothesis on $\mathcal{X}$, any sequence $(E_{i})_{i\geq 1}\subseteq\mathcal{E}_{\mathcal{H}}^{\vee}$ has a subsequence that converges point-wise on $\mathcal{X}$ to some $E\in\mathcal{E}_{\mathcal{H}}^{\vee}$.
Proof.
By Lemma 5 we can fix a minimal constraint $\Phi$ for $\mathcal{H}$. By Lemma 17, $\Lambda_{\Phi}$ is compact. For any $(E_{i})_{i\geq 1}\subseteq\mathcal{E}_{\mathcal{H}}^{\vee}$, we consider a corresponding sequence $(\lambda_{i})_{i\geq 1}\subseteq\Lambda_{\Phi}$, such that $E_{i}=1-\lambda_{i}\cdot\Phi$ for all $i\geq 1$. The existence of a subsequence of $(\lambda_{i})_{i\geq 1}$ converging to some $\lambda^{\star}\in\Lambda_{\Phi}$ implies that there is a subsequence of $(E_{i})_{i\geq 1}$ that converges point-wise to $1-\lambda^{\star}\cdot\Phi$, which is in $\mathcal{E}_{\mathcal{H}}^{\vee}$. ∎
Before proving the optimality of the dual e-class, we establish one last preliminary result.
Lemma 18.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$. Fix any countable set $S\subseteq\mathcal{X}$. Then, for each $E\in\mathcal{E}_{\mathcal{H}}$ we can find $E^{\prime}\in\mathcal{E}_{\mathcal{H}}^{\vee}$ such that $E^{\prime}(x)\geq E(x)$ for every $x\in S$.
Proof.
By Lemma 5 we can fix a minimal constraint $\Phi$ for $\mathcal{H}$. By Lemma 17, there is a finite matching set $S^{\prime}_{0}$ such that $\Lambda_{\Phi|S^{\prime}_{0}}$ is compact. Consider a sequence $(x_{i})_{i\geq 1}\subseteq\mathcal{X}$, dense in $\mathcal{X}$ and containing $S$. By Lemma 14, for all $i\geq 1$ we have a finite matching set $S^{\prime}_{i}$ containing $x_{i}$. For each $i\geq 1$, let $S_{i}=\bigcup_{j=0}^{i}S^{\prime}_{i}$. These sets are finite matching sets by Lemma 14. Let $D=\bigcup_{i=1}^{\infty}S_{i}$.
Fix $E\in\mathcal{E}_{\mathcal{H}}$. For all $i\geq 1$, $\mathcal{H}_{S_{i}}$ is a properly constrained on $S_{i}$ (Lemma 13). By Proposition 3 there is $\lambda_{i}\in\Lambda_{\Phi|S_{i}}$ such that $E_{\lambda_{i}}(x)\geq E(x)$, for all $x\in S_{i}$, where we let $E_{\lambda}=1-\lambda\cdot\Phi$ for $\lambda\in\mathbb{R}^{m}$. Since each $\lambda_{i}$ is in $\Lambda_{\Phi|S_{0}^{\prime}}$, which is compact by construction, $(\lambda_{i})_{i\geq 1}$ has a subsequence $(\lambda^{\star}_{i})_{i\geq 1}$ converging to some $\lambda^{\star}\in\Lambda_{\Phi|S_{0}^{\prime}}$. Note that $(S_{i})_{i\geq 1}$ is non-decreasing under set inclusion, namely $S_{i}\subseteq S_{i+1}$ for all $i\geq 1$. In particular, we can consider the subsequence $(S_{i}^{\star})_{i\geq 1}$ of $(S_{i})_{i\geq 1}$, obtained by selecting those indices that had been used to generate $(\lambda^{\star}_{i})_{i\geq 1}$, and still $D=\bigcup_{i=1}^{\infty}S^{\star}_{i}$. For any $i\geq 1$ and $j\geq i$, $\lambda^{\star}_{j}\in\Lambda_{\Phi|S^{\star}_{i}}$, so $\lambda^{\star}\in\bigcap_{i=1}^{\infty}\Lambda_{\Phi|S^{\star}_{i}}=\Lambda_{\Phi|D}$, by Lemma 15. Clearly, the convergence of $(\lambda^{\star}_{i})_{i\geq 1}$ to $\lambda^{\star}$ implies that $(\lambda_{i}^{\star}\cdot\Phi)_{i\geq 1}$ converges point-wise to $\lambda^{\star}\cdot\Phi$ on $\mathcal{X}$. Fix $k\geq 1$. For any $i\geq k$, $E_{\lambda^{\star}_{i}}(x)\geq E(x)$ if $x\in S^{\star}_{k}$, since $S^{\star}_{k}\subseteq S^{\star}_{i}$. So, for all $x\in S^{\star}_{k}$, $E_{\lambda^{\star}}(x)=\lim_{i\to\infty}E_{\lambda^{\star}_{i}}(x)\geq E(x)$. This holds for any $k\geq 1$, so $E_{\lambda^{\star}}(x)\geq E(x)$ for all $x\in S\subseteq D$.
To conclude we only need to show that $E_{\lambda^{\star}}\in\mathcal{E}_{\mathcal{H}}^{\vee}$. Since $\lambda^{\star}\in\Lambda_{\Phi|D}$, we have that $\lambda^{\star}\cdot\Phi(x)\leq 1$ for all $x\in D$. As $D\supseteq(x_{i})_{i\geq 1}$ is dense in $\mathcal{X}$ and $\lambda^{\star}\cdot\Phi$ is continuous, we deduce that $\lambda^{\star}\in\Lambda_{\Phi}$. ∎
Proposition 4.
Assume that $\mathcal{X}\subseteq\mathbb{R}^{n}$ is compact. Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$. Then, the optimal e-class exists and coincides with the dual e-class $\mathcal{E}_{\mathcal{H}}^{\vee}$.
Proof.
Fix $E\in\mathcal{E}_{\mathcal{H}}$. Only one of the following mutually contradictory statements is true.
  1. 1.
    There is $\varepsilon>0$ such that, for every $\hat{E}\in\mathcal{E}_{\mathcal{H}}^{\vee}$, there is $x\in\mathcal{X}$ such that $E(x)>\hat{E}(x)+\varepsilon$.
  2. 2.
    For every $\varepsilon>0$ there is $\hat{E}\in\mathcal{E}_{\mathcal{H}}^{\vee}$ such that, for all $x\in\mathcal{X}$, $E(x)\leq\hat{E}(x)+\varepsilon$.
Let us show that the claim (1) is false. First, note that since the space of continuous functions on a compact set in $\mathbb{R}^{n}$ is a separable metric space (under the uniform norm), there is a sequence $(\hat{E}_{i})_{i\geq 1}\subseteq\mathcal{E}_{\mathcal{H}}^{\vee}$ that is dense in $\mathcal{E}_{\mathcal{H}}^{\vee}$ (under the uniform convergence). If option (1) holds, there is $\varepsilon>0$ and a sequence $(x_{i})_{i\geq 1}\subseteq\mathcal{X}$ such that, for each $i\geq 1$, $E(x_{i})>\hat{E}_{i}(x_{i})+\varepsilon$. By Lemma 18, there is $\hat{E}\in\mathcal{E}_{\mathcal{H}}^{\vee}$ such that $\hat{E}(x_{i})\geq E(x_{i})>\hat{E}_{i}(x_{i})+\varepsilon$, for all $i\geq 1$. Since $(\hat{E}_{i})_{i\geq 1}$ is dense in $\mathcal{E}_{\mathcal{H}}^{\vee}$, there is an index $i^{\star}$ such that $\sup_{x\in\mathcal{X}}|\hat{E}(x)-\hat{E}_{i^{\star}}(x)|<\varepsilon$, a contradiction. So, (1) is false.
As the claim (2) must be true, we can find a sequence $(E_{i})_{i\geq 1}\subseteq\mathcal{E}_{\mathcal{H}}^{\vee}$, such that $E\preceq E_{i}+i^{-1}$ for all $i\geq 1$. By Corollary 2, there is a subsequence of $(E_{i})_{i\geq 1}$ that converges point-wise to some $E^{\star}\in\mathcal{E}_{\mathcal{H}}^{\vee}$. It follows that $E\preceq E^{\star}$, from which we deduce that $\mathcal{E}_{\mathcal{H}}^{\vee}$ is a majorising e-class. As all its elements are maximal (Lemma 10), it is the optimal e-class by Lemma 2. ∎
Theorem 1.
Let $\mathcal{H}$ be a properly constrained hypothesis. Then, the dual e-class $\mathcal{E}_{\mathcal{H}}^{\vee}$ is optimal.
Proof.
As all elements of $\mathcal{E}_{\mathcal{H}}^{\vee}$ are maximal (Lemma 10), by Lemma 2 it is enough to show that $\mathcal{E}_{\mathcal{H}}^{\vee}$ is a majorising e-class. Fix $E\in\mathcal{E}_{\mathcal{H}}$. By Lemma 5 we can fix a minimal constraint $\Phi$ for $\mathcal{H}$. By Lemma 17, there is a finite matching set $S^{\star}\subseteq\mathcal{X}$ such that any closed set $S\subseteq\mathcal{X}$ containing $S^{\star}$ is a matching set that induces a compact $\Lambda_{\Phi|S}$. Since $\mathcal{X}$ is closed, it is sigma-compact, and we can find a sequence $(K_{i})_{i\geq 1}$ of compact matching sets, such that $S^{\star}\subseteq K_{i}\subseteq K_{i+1}$, for all $i\geq 1$, and $\mathcal{X}=\bigcup_{i=1}^{\infty}K_{i}$. By construction, $\mathcal{H}_{K_{i}}$ is properly constrained (Lemma 13) and $\Phi\big{|}_{K_{i}}$ a constraint for it (Lemma 12). From Proposition 4, for each $i\geq 1$ there is $\lambda_{i}\in\Lambda_{\Phi|K_{i}}$ such that $E_{\lambda_{i}}(x)\geq E(x)$ for all $x\in K_{i}$, where as usual $E_{\lambda}=1-\lambda\cdot\Phi$. For all $i\geq 1$, because $K_{i}\subseteq K_{i+1}$ we have $\Lambda_{\Phi|K_{i}}\supseteq\Lambda_{\Phi|K_{i+1}}$. In particular, for any $j\geq 1$, $(\lambda_{i})_{i\geq j}$ is contained in $\Lambda_{\Phi|K_{j}}$, which is compact by Lemma 17. So, there is a subsequence $(\lambda^{\prime}_{i})_{i\geq 1}$ of $(\lambda_{i})_{i\geq 1}$ that converges to $\lambda^{\star}\in\bigcap_{j\geq 1}\Lambda_{\Phi|K_{j}}$. Let $(K^{\prime}_{i})_{i\geq 1}$ be the corresponding subsequence of $(K_{i})_{i\geq 1}$, obtained selecting the same indices. By Lemma 15, $\lambda^{\star}\in\Lambda_{\Phi}$, so $E_{\lambda^{\star}}\in\mathcal{E}_{\mathcal{H}}^{\vee}$. Fix any $x\in\mathcal{X}$. There is an index $i^{\star}\geq 1$ such that $x\in K_{i}^{\prime}$ for all $i\geq i^{\star}$. Thus, $E_{\lambda^{\prime}_{i}}(x)\geq E(x)$ for all $i\geq i^{\star}$, and so $E_{\lambda^{\star}}(x)\geq E(x)$. As the choice of $x$ has been arbitrary, $E_{\lambda^{\star}}\succeq E$. We conclude that $\mathcal{E}_{\mathcal{H}}^{\vee}$ is a majorising e-class. ∎
9 Extensions
9.1 Testing finitely (non-properly) constrained hypotheses
So far we have mostly focused on those finitely constrained hypotheses $\mathcal{H}$ that are also properly constrained. Indeed, if $\mathcal{H}$ is not properly constrained, by Lemma 8 and Corollary 1 the optimal e-class does not exist. However, we can still characterise an “optimal” e-variable testing procedures. The key idea is the following result, whose proof is detailed in Appendix A.
Lemma 19.
Let $\mathcal{H}$ be a finitely constrained hypothesis on $\mathcal{X}$. Let $\mathcal{X}_{0}=\bigcup_{P\in\mathcal{H}}\operatorname*{Supp}P$. Then, $\mathcal{X}_{0}$ is closed and $\mathcal{H}$ is a properly constrained hypothesis on $\mathcal{X}_{0}$.
Lemma 19 tells us that any finitely hypothesis $\mathcal{H}$ on $\mathcal{X}$ can be seen as a properly constrained hypothesis on some smaller set $\mathcal{X}_{0}$. By the definition of $\mathcal{X}_{0}$, we know that observing $x_{t}\notin\mathcal{X}_{0}$ is incompatible with the hypothesis $\mathcal{H}$, as such $x_{t}$ is not in the support of any $P\in\mathcal{H}$. So, we can design the following testing procedure. Assuming that $\mathcal{H}$ holds, we see it as a hypothesis on $\mathcal{X}_{0}$ and we design a testing game on $\mathcal{X}_{0}$ (as in Definition 2) restricted to the optimal e-class of $\mathcal{H}$ on $\mathcal{X}_{0}$. At each new observation $x_{t}$, if $x_{t}\notin\mathcal{X}_{0}$ we automatically reject $\mathcal{H}$. Conversely, if $x_{t}\in\mathcal{X}_{0}$, we let the player in the testing game observe $x_{t}$, and as usual we reject $\mathcal{H}$ if the cumulative reward exceeds the threshold $\log(1/\delta)$, with $\delta\in(0,1)$ the type II confidence level.
9.2 Loosely constrained hypotheses
Finitely constrained hypotheses involve the equality $\langle P,\Phi\rangle=0$. Here, we relax this condition.
Definition 10 (Loosely constrained hypotheses).
A hypothesis $\mathcal{H}$ on $\mathcal{X}$ is loosely constrained if there are two continuous functions $\Phi^{\prime}:\mathcal{X}\to\mathbb{R}^{m^{\prime}}$ and $\Phi^{\prime\prime}:\mathcal{X}\to\mathbb{R}^{m^{\prime\prime}}$ such that, denoted as $\Phi:\mathcal{X}\to\mathbb{R}^{m^{\prime}+m^{\prime\prime}}$ the mapping $x\mapsto\Phi(x)=(\Phi^{\prime}(x),\Phi^{\prime\prime}(x))$, we have that $0\in\operatorname*{conv}\Phi(\mathcal{X})$ and
=$\mathcal{H}=\{P\in\mathcal{P}_{\Phi^{\prime}}\cap\mathcal{P}_{\Phi^{\prime\prime}}\;:\;\langle P,\Phi^{\prime}\rangle=0\,,\;\langle P,\Phi^{\prime\prime}\rangle\geq 0\}\,,$
where the inequality holds component-wise. We call $\Phi^{\prime}$ a tight constraint and $\Phi^{\prime\prime}$ a slack constraint. If moreover $0\in\operatorname*{rel\,int}(\operatorname*{conv}\Phi(\mathcal{X}))$, $\mathcal{H}$ is said to be loosely properly constrained.
For any properly loosely constrained hypothesis the optimal e-class exists. This is made explicit in the next proposition (see Appendix A for the proof).
Proposition 5.
Let $\mathcal{H}$ be a properly loosely constrained hypothesis on $\mathcal{X}$, with $\Phi^{\prime}:\mathcal{X}\to\mathbb{R}^{m^{\prime}}$ and $\Phi^{\prime\prime}:\mathcal{X}\to\mathbb{R}^{m^{\prime\prime}}$ tight and slack constraints. Define
[…]
Then, the optimal e-class for $\mathcal{H}$ exists and coincides with
~=$\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}=\big{\{}1-\lambda^{\prime}\cdot\Phi^{\prime}-\lambda^{\prime\prime}\cdot\Phi^{\prime\prime}\;:\;(\lambda^{\prime},\lambda^{\prime\prime})\in\tilde{\Lambda}_{\Phi^{\prime},\Phi^{\prime\prime}}\big{\}}\,.$
We remark that the set =$\mathcal{H}=\{P\in\mathcal{P}_{\Phi}\,:\,\langle P,\Phi\rangle=0\}$, where $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ is a continuous function, is non-empty (and thus a hypothesis) if, and only if, $0\in\operatorname*{conv}\Phi(\mathcal{X})$ (by Lemma 22 and Lemma 24). Conversely, when we allow for inequality constraints, as in the definition of loosely constrained hypotheses, this is not anymore a necessary condition. More explicitly, =$\mathcal{H}=\{P\in\mathcal{P}_{\Phi^{\prime}}\cap\mathcal{P}_{\Phi^{\prime\prime}}\,:\,\langle P,\Phi^{\prime}\rangle=0\,,\,\langle P,\Phi^{\prime\prime}\rangle\geq 0\}$ can be non empty if $0\notin\operatorname*{conv}\Phi(\mathcal{X})$, (where $\Phi^{\prime}$ and $\Phi^{\prime\prime}$ are continuous functions on $\mathcal{X}$ and Φ=($\Phi=(\Phi^{\prime},\Phi^{\prime\prime})$). However, here we only focus on the case $0\in\operatorname*{conv}\Phi(\mathcal{X})$, as required by Definition 10, and defer to future work the analysis of the setting of generic hypotheses defined via loose constraints. We also note that Lemma 19 cannot be trivially extended to the case of loosely constrained hypotheses, and it is not always possible to restrict the domain $\mathcal{X}$ of a loosely constrained hypothesis to make it properly loosely constrained. For an example where this is not possible, consider the case of a loosely constrained hypothesis on $\mathcal{X}=\mathbb{R}$, with tight constraint $\Phi^{\prime}:x\mapsto 1-x$ and slack constraint $\Phi^{\prime\prime}:x\mapsto x^{2}-1$.
10 Algorithmic mean estimation
In Clerico (2024), the notion of optimal e-class was developed in the context of applying the testing framework to mean estimation for random variables bounded in $[0,1]$. The class of constrained hypotheses that we have introduced allows for a substantial extension of those results.
10.1 Mean estimation for bounded random variables
Let $\mathcal{X}$ be a compact set in $\mathbb{R}^{n}$ and let $\mathcal{C}$ denote its convex hull. We observe independent draws $(X_{i})_{i\geq 1}$ from an unknown distribution $P^{\star}\in\mathcal{P}_{\mathcal{X}}$, whose mean $\mu^{\star}\in\mathcal{C}$ needs to be estimated. It is possible to generate an anytime-valid sequence of confidence sets (usually referred to as confidence sequence) via sequential hypothesis testing (see, e.g., Ramdas and Wang, 2024). More precisely, we are looking for a sequence of random sets $(S_{t})_{t\geq 1}$, adapted to $(X_{i})_{i\geq 1}$ and such that
$\mathbb{P}\big{(}\mu^{\star}\in S_{t}\,,\;\forall t\geq 1\big{)}\geq 1-\delta\,,$
where $\mathbb{P}$ denotes the law of the sequence of observations and $\delta\in(0,1)$ is a fixed confidence level.
Let $\mathcal{P}_{1}$ denote the set of measures $P\in\mathcal{P}_{\mathcal{X}}$ satisfying $\langle P,\|X\|_{1}\rangle<+\infty$. Clearly, for $P\in\mathcal{P}_{1}$ the mean $\langle P,X\rangle$ is well defined. For $\mu\in\mathcal{C}$, let
$\mathcal{H}_{\mu}=\left\{P\in\mathcal{P}_{1}\;:\;\langle P,X\rangle=\mu\right\}\,.$
Clearly, $\mathcal{H}_{\mu}$ is a hypothesis on $\mathcal{X}$ (it is non-empty by Lemma 22) and it is finitely constrained, with $\Phi_{\mu}:x\mapsto\mu-x$ a constraint (as $0\in\operatorname*{conv}\Phi_{\mu}(\mathcal{X})$ because $\mu\in\mathcal{C}$). Moreover, if $\mu\in\operatorname*{rel\,int}\mathcal{C}$, $\mathcal{H}_{\mu}$ is a properly constrained hypothesis on $\mathcal{X}$. For this reason, for the sake of simplicity here we assume as known that $\mu^{\star}\in\operatorname*{rel\,int}\mathcal{C}$. Therefore, we are only interested in confidence sets $S_{t}\subseteq\operatorname*{rel\,int}\mathcal{C}$. We note, however, that the general case can be addressed by dealing with the values of $\mu$ on the relative boundary of $\mathcal{C}$ as discussed in Section 9.1. A valid confidence sequence for $\mu^{\star}\in\operatorname*{rel\,int}\mathcal{C}$ with confidence level $\delta\in(0,1)$ is $(S_{t})_{t\geq 1}$, where
$S_{t}=\big{\{}\mu\in\operatorname*{rel\,int}\mathcal{C}\;:\;R_{t}(\mu)\leq\log(1/\delta)\big{\}}\,.$
Here, $R_{t}(\mu)$ denotes the cumulative reward, up to round $t$, of a player who is playing a testing game as in Definition 2 on the null hypothesis $P^{\star}\in\mathcal{H}_{\mu}$, observing the sequence $(X_{i})_{i\geq 1}$. From the fact that $\{\mu^{\star}=S_{t}\,,\,\forall t\geq 1\}\iff\{R_{t}(\mu^{\star})\leq\log(1/\delta)\,,\,\forall t\geq 1\}$, which happens with probability at least $1-\delta$ by Lemma 1, it follows that $(S_{t})_{t\geq 1}$ is a confidence sequence.
Of course, in order to get a tight confidence sequence, we shall restrict the testing games to the optimal e-classes (see also the discussion in Clerico, 2024). For every $\mu\in\operatorname*{rel\,int}\mathcal{C}$, $\mathcal{H}_{\mu}$ is a properly constrained hypothesis and we can leverage the theory that we have developed so far. In particular, we have $\Lambda_{\Phi_{\mu}}=\{\lambda\in\mathbb{R}^{n}\,:\,\sup_{x\in\mathcal{X}}\lambda\cdot(\mu-x)\leq 1\}$, and the optimal e-class is
$\mathcal{E}_{\mathcal{H}_{\mu}}^{\vee}=\left\{x\mapsto 1+\lambda\cdot(x-\mu)\,:\,\lambda\in\Lambda_{\Phi_{\mu}}\right\}\,.$
When $\mathcal{X}=[0,1]$, $\mathcal{E}_{\mathcal{H}_{\mu}}^{\vee}$ is the coin-betting e-class studied in Clerico (2024), which had been used for mean estimation by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023).
10.2 Mean estimation for heavy-tailed distributions
Consider the case $\mathcal{X}=\mathbb{R}$. We want to estimate the mean $\mu^{\star}$ of an unknown probability distribution $P^{\star}\in\mathcal{P}_{\mathbb{R}}$, under the assumption that $\langle P^{\star},X^{2}\rangle\leq 1$. Note that this is a heavy-tail mean estimation problem, as we only have an upper bound on a moment of $P^{\star}$. Let $(X_{i})_{i\geq 1}$ be a sequence of independent draws from $P^{\star}$. As we have just done in Section 10.1, we aim at obtaining a confidence sequence $(S_{t})_{t\geq 1}$ for $\mu^{\star}$. Clearly, the fact that $\langle P^{\star},X^{2}\rangle\leq 1$ implies that $\mu^{\star}\in[-1,1]$. For $\mu\in[-1,1]$, let
$\mathcal{H}_{\mu}=\big{\{}P\in\mathcal{P}_{\mathbb{R}}\;:\;\langle P,X^{2}\rangle\leq 1\,,\;\langle P,X\rangle=\mu\big{\}}\,,$
where we note that the fact that $\langle P,X^{2}\rangle\leq 1$ automatically implies that $\langle P,X\rangle$ is well defined. We remark that $\mathcal{H}_{1}=\{\delta_{1}\}$ and $\mathcal{H}_{-1}=\{\delta_{-1}\}$. On the other hand, it is easy to check that, for $\mu\in(-1,1)$, $\mathcal{H}_{\mu}$ is a loosely properly constrained hypothesis on $\mathbb{R}$, where $\Phi^{\prime}_{\mu}:x\mapsto\mu-x$ is a tight constraint and $\Phi^{\prime\prime}:x\mapsto 1-x^{2}$ is a slack constraint. Denoting as $R_{t}(\mu)$ the cumulative reward (up to round $t$) of a player betting in a testing game on $\mathcal{H}_{\mu}$ and observing the sequence $(X_{i})_{i\geq 1}$, with the same reasoning as in Section 10.1, we deduce that
$S_{t}=\big{\{}\mu\in(-1,1)\;:\;R_{t}(\mu)\leq\log(1/\delta)\big{\}}\cup U_{t}$
defines a confidence sequence for $\mu^{\star}$, where $U_{t}=\{1\}$ if for every $s\leq t$ all $X_{s}=1$, $U_{t}=\{-1\}$ if for every $s\leq t$ all $X_{s}=-1$, and $U_{t}=\varnothing$ otherwise.
Once more, we shall restrict the testing games to the optimal e-classes. By Proposition 5, for $\mu\in(-1,1)$ the optimal e-class is
~$\tilde{\mathcal{E}}_{\mathcal{H}_{\mu}}^{\vee}=\big{\{}x\mapsto 1+\alpha(x-\mu)+\beta(x^{2}-1)\;:\;(\alpha,\beta)\in\tilde{\Lambda}_{\Phi^{\prime}_{\mu},\Phi^{\prime\prime}}\}\,,$
where $\tilde{\Lambda}_{\Phi^{\prime}_{\mu},\Phi^{\prime\prime}}=\big{\{}(\alpha,\beta)\in\mathbb{R}^{2}\,:\,\alpha^{2}+4\mu\alpha\beta+4\beta^{2}-4\beta\leq 0\}$ is the convex hull of an ellipse lying on the half-plane where $\beta\geq 0$.
[…]
We note that the problem of finding confidence sequences for heavy-tailed random variables via betting strategies has been addressed in the literature. Agrawal et al. (2021) obtained anytime-valid lower bounds for real random variables with bounded $(1+\varepsilon)$-th (non-centred) moment, via an approach that implicitly uses e-variables, as their Lemma 17 is in fact a regret upper bound for the cumulative reward of a player in a testing game as in Definition 2. Notably, our results show that their analysis considered testing games restricted to the optimal e-class. On the other hand, Wang and Ramdas (2023) leveraged ideas from Catoni (2012) to set up testing games in order to find confidence sequences for the mean of real random variables with bounded $(1+\varepsilon)$-th central moment. However, the theory that we have developed in this work shows that their choice of restriction for the e-variables does not correspond to the optimal e-class.
11 Conclusion
This work helps to deepen the understanding of the notion of optimal e-class introduced in Clerico (2024) for hypothesis testing via e-variables, focusing on its characterisation for non-parametric hypotheses defined by a finite number of continuous constraints. We now conclude with some final remarks and possible directions for future research.
First, it is worth noticing that key concepts of maximal e-variable, and majorising and optimal e-class that we discuss in this work are strongly connected to the classical statistical notions of admissibility and completeness (see, e.g., Section 1.8 in Lehmann and Romano, 2022). Indeed, with this terminology, a maximal e-variable can be thought of as an admissible e-variable, in the sense that no other e-variable uniformly dominates it, a majorising e-class is a complete class of e-variables, and the optimal e-class represents the minimal complete class. Notably, the idea of admissibility has already been explored in the e-variable literature. Indeed, Ramdas et al. (2022) considered a notion of admissible wealth processes, which in some contexts is closely related to the maximality for e-variables discussed here. However, their results do not suffice to characterise the optimal e-class in the setting that we study, as already pointed out by Clerico (2024). Admissibility has also been recently examined in a different context involving e-variables, in the problem of “merging” e-variables (Wang, 2024). Exploring the connections between those ideas and the concepts discussed in this work could be a promising direction for future research.
To establish our results, we employed a standard functional analytic approach. We began by proving the optimality of the dual e-class for a set $\mathcal{X}$ with finite cardinality, then extended it to the case of a compact $\mathcal{X}$ using a dense countable subset, and finally leveraged sigma-compactness for a general closed $\mathcal{X}$. Our proofs are self-contained and avoid relying on advanced tools from functional analysis or measure theory. We mention, however, that an alternative approach, leveraging bipolar-like theorems, might potentially yield similar conclusions, though it remains unclear at this stage whether it would achieve the desired results and simplify the analysis. Connections between e-variables and polar sets have already been exploited in Larsson et al. (2024), building upon recent technical advances in characterising the bipolar of subsets of probability measures (see, e.g., Langner and Svindland (2022) and the references therein). Exploring the potential application of this approach to our framework is an interesting direction for future work, which might also help addressing more general settings.
Finally, open questions remain regarding the minimal conditions, on the hypothesis $\mathcal{H}$, necessary for the existence of the optimal e-class, and the extent to which such e-class can be explicitly characterised, when it does exist. A simpler, perhaps more approachable, problem is whether it is possible to extend our analysis to hypotheses defined by countably many scalar constraints, which would allow to determine the optimal e-classes to estimate the mean of random variables with a known bound on the moment generating function (e.g., sub-Gaussian random variables).
Acknowledgements
I would like to thank Gergely Neu, Peter Grünwald, Gabor Lugosi, Nishant Mehta, Hamish Flynn, Ludovic Schwartz, Claudia Chanu, and Riccardo Camerlo for the insightful discussions that contributed to this work, and Nick Koning for pointing out interesting connections with related literature. This project was funded by the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (grant agreement 950180).
Appendix A Omitted proofs
Lemma 4.
If $\mathcal{X}$ is countable and $\sup_{P\in\mathcal{H}}P(\{x\})>0$ for all $x\in\mathcal{X}$, the optimal e-class exists.
Proof.
By Lemma 2, it is enough to show that every e-variable is majorised by a maximal e-variable, which implies that the set of all maximal e-variables is a majorising class. Fix $E_{0}\in\mathcal{E}_{\mathcal{H}}$. If $E_{0}$ is already maximal, we are done. So, we focus on the case of an $E_{0}$ that is not maximal.
Consider the transfinite sequence, starting from $E_{0}$ and indexed in $\omega_{1}$ (the smallest uncountable ordinal), defined as follows. For any countable successor ordinal $\beta\geq 1$, let $E_{\beta}=E_{\beta-1}$ if $E_{\beta-1}$ is maximal, otherwise pick any $E\in\mathcal{E}_{\mathcal{H}}$ such that $E\succ E_{\beta-1}$ and let $E_{\beta}=E$. For any limiting ordinal $\gamma\in\omega_{1}$, let $(\beta_{n})_{n\geq 1}\subseteq\omega_{1}$ be an increasing sequence of ordinals converging to $\gamma$ (with respect to the order topology). For $x\in\mathcal{X}$, let $A(x)=1/\sup_{P\in\mathcal{H}}P(\{x\})\in[1,+\infty)$. Note that for any $E\in\mathcal{E}_{\mathcal{H}}$ and any $x\in\mathcal{X}$, $E(x)=A(x)E(x)\sup_{P\in\mathcal{H}}P(\{x\})\leq A(x)\sup_{P\in\mathcal{H}}\langle P,E\rangle\leq A(x)$. So, for each $x\in\mathcal{X}$, the sequence $(E_{\beta_{n}}(x))_{n\geq 1}$ is non-decreasing and bounded in $[0,A(x)]$, and hence admits a finite limit. Therefore, we can define the point-wise limit $E_{\gamma}=\lim_{n\to\infty}E_{\beta_{n}}$. Moreover, by Beppo Levi theorem, $E_{\gamma}$ is Borel and $\langle P,E_{\gamma}\rangle\leq 1$, for any $P\in\mathcal{H}$. So, $E_{\gamma}\in\mathcal{E}_{\mathcal{H}}$.
Now, consider the set $B\subseteq\omega_{1}$ of the indices where the transfinite sequence has jumps (namely, $E_{\beta}\succ E_{\beta-1}$ if, and only if, $\beta\in B$). Since $E_{0}$ is not maximal, $B$ is non-empty. For each $\beta\in B$ we can select a pair $(x_{\beta},q_{\beta})\in\mathcal{X}\times\mathbb{Q}$ such that $E_{\beta-1}(x_{\beta})<q_{\beta}<E_{\beta}(x_{\beta})$. This defines an injection $B\to\mathcal{X}\times\mathbb{Q}$, whose existence implies that $B$ is countable, because $\mathcal{X}\times\mathbb{Q}$ is countable. Thus, $\beta^{\star}=\sup B\in\omega_{1}$, as it is the supremum of a countable subset of $\omega_{1}$. By construction, $E_{\beta^{\star}}\succeq E_{0}$ and $E_{\beta^{\star}}$ is maximal, since it must be that $E_{\beta^{\star}+1}=E_{\beta^{\star}}$, as otherwise $\beta^{\star}+1$ would be in $B$. ∎
Lemma 5.
Every properly constrained hypothesis admits a minimal constraint.
Proof.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$ and $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ be a proper constraint. Denote as $U$ the affine hull $\operatorname*{aff}\Phi(\mathcal{X})$, and let $\Pi_{U}$ be the orthogonal projection from $\mathbb{R}^{m}$ to $U$. Let $\Phi_{U}=\Pi_{U}\circ\Phi$ and $\mathcal{H}_{U}=\{P\in\mathcal{P}_{\Phi_{U}}\,:\,\langle P,\Phi_{U}\rangle=0\}$. Since $\Phi(\mathcal{X})\subseteq U$, we have that $\Phi_{U}(x)=\Phi(x)$ for all $x\in\mathcal{X}$. It follows immediately that $\mathcal{H}=\mathcal{H}_{U}$. Using that $U\cong\mathbb{R}^{m^{\prime}}$ (for some $m^{\prime}\leq m$), we can define a linear isomorphism $I:U\to\mathbb{R}^{m^{\prime}}$. Then, $\langle P,\Phi_{U}\rangle=0$ if, and only if, $\langle P,I\circ\Phi_{U}\rangle=0$. It follows that $I\circ\Phi_{U}$ is a proper constraint. By construction, $0\in\operatorname*{int}(\operatorname*{conv}(I\circ\Phi_{U}(\mathcal{X})))$, so $I\circ\Phi_{U}$ is a minimal constraint. ∎
Lemma 6.
Let $\mathcal{H}$ be a properly constrained hypothesis and $\Phi$ a proper constraint. $\Phi$ is minimal if, and only if, all its components are linearly independent scalar functions.
Proof.
Let $\dim(\operatorname*{Span}\Phi)=m^{\prime}$. By Lemma 23, $m^{\prime}$ is also the dimension of the affine hull of $\Phi(\mathcal{X})$. Hence, $\operatorname*{aff}\Phi(\mathcal{X})=\mathbb{R}^{m}$ if, and only if, $m^{\prime}=m$. If this is the case, then the interior and relative interior of $\operatorname*{conv}\Phi(\mathcal{X})$ coincide, and so $0\in\operatorname*{rel\,int}(\operatorname*{conv}\Phi(\mathcal{X}))$ implies that $0\in\operatorname*{int}(\operatorname*{conv}\Phi(\mathcal{X}))$. On the other hand, if $m^{\prime}<m$ the interior of $\operatorname*{conv}\Phi(\mathcal{X})$ is empty, and $\Phi$ cannot be minimal. ∎
Lemma 7.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$. Then, for every $x\in\mathcal{X}$ there exists a $P\in\mathcal{H}$ whose support has finite cardinality and such that $x\in\operatorname*{Supp}P$.
Proof.
Let $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ be a proper constraint and denote as $\mathcal{Y}$ its image. Now, fix $x_{0}\in\mathcal{X}$ and let $y_{0}=\Phi(x_{0})$. By Corollary 3, since $0\in\operatorname*{rel\,int}(\operatorname*{conv}\mathcal{Y})$, we can find $P^{\prime}\in\mathcal{P}_{\mathcal{Y}}$, with zero mean and such that $P^{\prime}(\{y_{0}\})>0$, and whose support has finitely many (say $r+1$) elements. So, for some $r\geq 0$ we can write $\operatorname*{Supp}P^{\prime}=\{y_{0},y_{1},\dots,y_{r}\}$. We can choose $r$ points $\{x_{1},\dots,x_{r}\}\subseteq\mathcal{X}$ such that $\Phi(x_{i})=y_{i}$ for any $i=1\dots r$. Then, we define $P=\sum_{i=0}^{r}P^{\prime}(\{y_{i}\})\delta_{x_{i}}$. Clearly, $\operatorname*{Supp}P$ has finitely many elements (so, $\langle P,\|\Phi\|_{1}\rangle<+\infty$) and $P(\{x_{0}\})=P^{\prime}(\{y_{0}\})>0$. Moreover, $\langle P,\Phi\rangle=\langle P^{\prime},Y\rangle=0$ by construction, so $P\in\mathcal{H}$. ∎
Lemma 8.
Let $\mathcal{H}$ be a finitely constrained hypothesis on $\mathcal{X}$. $\mathcal{H}$ is properly constrained if, and only if, $\sup_{P\in\mathcal{H}}P(\{x\})>0$ for all $x\in\mathcal{X}$. Moreover, if $\mathcal{H}$ is properly constrained, every constraint $\Phi$ for $\mathcal{H}$ is a proper constraint and there is $A:\mathcal{X}\to[1,+\infty)$ such that $E\preceq A$ for any $E\in\mathcal{E}_{\mathcal{H}}$.
Proof.
Let $\Phi$ be a constraint for $\mathcal{H}$. Note that if $\Phi$ is identically null, then $\mathcal{H}=\mathcal{P}_{\mathcal{X}}$ and $\Phi$ is a proper constraint, as $\operatorname*{rel\,int}\{0\}=\{0\}$. So, we only need to consider the case where $\Phi$ is not identically null.
First, if $\mathcal{H}$ is properly constrained, by Lemma 7, $\sup_{P\in\mathcal{H}}P(\{x\})>0$, for all $x\in\mathcal{X}$.
Now, assume that $\sup_{P\in\mathcal{H}}P(\{x\})>0$ for all $x\in\mathcal{X}$, and let $\mathcal{H}$ be a finitely constrained hypothesis. Let $\Phi$ be a constraint. We show that $\Phi$ is a proper constraint by contradiction. Denote as $\mathcal{C}$ the convex hull of $\Phi(\mathcal{X})$, and let $\partial\mathcal{C}$ be its relative boundary. Assume that $\Phi$ is not a proper constraint. Then, $0\in\partial\mathcal{C}$. Since $\mathcal{H}$ is non-empty and $\Phi$ is not identically null, there are at least two distinct points in $\Phi(\mathcal{X})$. Hence, $\mathcal{C}$ has non-empty relative interior and there is $x_{0}\in\mathcal{X}$ such that $\Phi(x_{0})\in\operatorname*{rel\,int}\mathcal{C}$. By assumption, there is $P\in\mathcal{H}$ such that $P(\{x_{0}\})>0$, so $x_{0}\in\operatorname*{Supp}P$. But by Lemma 24, $\operatorname*{Supp}P\subseteq\Phi^{-1}(\partial\mathcal{C})$ for each $P\in\mathcal{H}$. This is a contradiction since $x_{0}\in\Phi^{-1}(\operatorname*{rel\,int}\mathcal{C})$, which is an open set [^4] Indeed, we can see $\Phi$ as a continuous function from $\mathcal{X}$ to $\operatorname*{aff}\Phi(\mathcal{X})$, and $\operatorname*{rel\,int}\mathcal{C}$ is an open set of $\operatorname*{aff}\Phi(\mathcal{X})$, so its counter-image is open. that has an empty intersection with $\Phi^{-1}(\partial\mathcal{C})$.
Note that what shown so far also implies that any constraint of a properly constrained hypothesis is a proper constraint.
Finally, let $\mathcal{H}$ be properly constrained and $A:x\mapsto 1/\sup_{P\in\mathcal{H}}P(\{x\})$, which defines a function $\mathcal{X}\to[1,+\infty)$. Then, for any e-variable $E\in\mathcal{H}$ and any $x\in\mathcal{X}$, we have
$E(x)\leq A(x)E(x)\sup_{P\in\mathcal{H}}P(\{x\})\leq A(x)\sup_{P\in\mathcal{H}}\langle P,E\rangle\leq A(x)\,,$
and we conclude. ∎
Lemma 9.
Let $\mathcal{H}$ be a finitely constrained hypothesis on $\mathcal{X}$. $\mathcal{H}$ is properly constrained if, and only if, there is $P\in\mathcal{H}$ such that $\operatorname*{Supp}P=\mathcal{X}$.
Proof.
Assume that $\mathcal{H}$ is properly constrained. Fix a sequence $(x_{i})_{i\geq 1}\subseteq\mathcal{X}$, dense in $\mathcal{X}$. For each $i\geq 1$, by Lemma 8 there is a $P_{i}\in\mathcal{H}$ such that $P_{i}(\{x\})>0$. Let $\Phi$ be a constraint. Since $P_{i}\in\mathcal{H}$ we have $P\in\mathcal{P}_{\Phi}$, and so $\zeta_{i}=\langle P_{i},\|\Phi\|_{1}\rangle$ is finite. Let $\alpha_{i}=2^{-i}/(1+\zeta_{i})$. Then, we have $\xi=\sum_{i=1}^{\infty}\alpha_{i}\in(0,1]$. Let $P=\sum_{i=1}^{\infty}\frac{\alpha_{i}}{\xi}P_{i}$. By construction, $P\in\mathcal{P}_{\mathcal{X}}$. Moreover, $\langle P,\|\Phi\|_{1}\rangle=\xi\sum_{i=1}^{\infty}\alpha_{i}z_{i}\leq\xi\sum_{i}^{\infty}2^{-i}=\xi$, and $\langle P,\Phi\rangle=0$. It follows that $P\in\mathcal{H}$. As, $\operatorname*{Supp}\Phi\supseteq(x_{i})_{i\geq 1}$, which is dense, we conclude that $\operatorname*{Supp}P=\mathcal{X}$.
Conversely, let $\mathcal{H}$ be finitely constrained and not properly constrained. For every constraint $\Phi$, it must be that $0\in\mathcal{C}=\operatorname*{conv}\Phi(\mathcal{X})$ (see Lemma 24). Since there is no proper constraint, 00 must be in the relative boundary $\partial\mathcal{C}$ of $\mathcal{C}$. Again by Lemma 24, any $P\in\mathcal{H}$ is supported on $\Phi^{-1}(\partial\mathcal{C})$. As the relative interior of a non-empty set is non-empty, $\operatorname*{Supp}P\subseteq\Phi^{-1}(\partial\mathcal{C})\subsetneq\mathcal{X}$. ∎
Lemma 12.
Let $\mathcal{H}$ be a finitely constrained hypothesis on $\mathcal{X}$ and $\Phi$ a constraint. A closed set $S\subseteq\mathcal{X}$ is compatible if, and only if, $0\in\operatorname*{conv}\Phi(S)$. In such case, $\mathcal{H}_{S}$ is a finitely constrained hypothesis on $S$ and $\Phi\big{|}_{S}$ is a constraint for it. In particular, $\mathcal{H}_{S}=\{P\in\mathcal{P}_{\Phi|S}\,:\,\langle P,\Phi\big{|}_{S}\rangle=0\}$.
Proof.
If $0\in\operatorname*{conv}\Phi(S)$, Lemma 22 implies that there is a finite set $\{x_{1},\dots,x_{r}\}\subseteq S$ such that $\sum_{i=1}^{r}\alpha_{i}\Phi(x_{i})=0$, with the $\alpha_{i}$ non-negative coefficients that sum to $1$. In particular, $P=\sum_{i=1}^{r}\alpha_{i}\delta_{x_{i}}$ is in $\mathcal{H}_{S}$, and so $S$ is compatible. On the other hand, if $S$ is compatible, then $0\in\operatorname*{conv}\Phi(S)$ by Lemma 24.
Now, let $\Phi$ be a constraint for $\mathcal{H}$. Then $\Phi\big{|}_{S}:S\to\mathbb{R}^{m}$ is continuous, as it is the restriction of a continuous function. Let $P\in\mathcal{H}_{S}$. Then, $P\in\mathcal{P}_{\Phi}\cap\mathcal{P}_{S}=\mathcal{P}_{\Phi|S}$, and since $\mathcal{H}_{S}\subseteq\mathcal{H}$, we have that $\langle P,\Phi\big{|}_{S}\rangle=\langle P,\Phi\rangle=0$. So, $\mathcal{H}_{S}\subseteq\{P\in\mathcal{P}_{\Phi|S}\;:\;\langle P,\Phi\big{|}_{S}\rangle=0\}$. Now, if $P\in\mathcal{P}_{\Phi|S}$ and $\langle P,\Phi\big{|}_{S}\rangle=0$, then $\langle P,\Phi\rangle=0$. So $P\in\mathcal{H}$ and, since $P\in\mathcal{P}_{S}$, we also have $P\in\mathcal{H}_{S}$. ∎
Lemma 13.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$ and $S\subseteq\mathcal{X}$ a compatible set. $S$ is a matching set if, and only if, $\mathcal{H}_{S}$ is a properly constrained hypothesis on $S$.
Proof.
This is a direct consequence of Lemma 9. ∎
Lemma 14.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$. A finite union of matching sets is a matching set. In particular, every finite subset of $\mathcal{X}$ is contained in a finite matching set.
Proof.
Let $\{S_{1},\dots,S_{N}\}$ be a finite family of matching sets. For each $S_{i}$ there is $P_{i}\in\mathcal{H}$ such that $\operatorname*{Supp}P_{i}=S_{i}$. Let $P=N^{-1}\sum_{i=1}^{N}P_{i}$. Then, $P\in\mathcal{H}$ by convexity, so $S^{\prime}=\operatorname*{Supp}P$ is a matching set. Moreover, $S^{\prime}$ is the closure of $\bigcup_{i=1}^{N}S_{i}$, but since every $S_{i}$ is closed we have $S^{\prime}=\bigcup_{i=1}^{m}S_{i}$. Now, let $S\subseteq\mathcal{X}$ be any finite set. By Lemma 7, for every $x\in S$ there is a $P_{x}\in\mathcal{H}$ with finite support $S_{x}$ that contains $x$. Clearly, $\bigcup_{x\in S}S_{x}$ contains $S$, and it is a matching set from what we have just shown, as it is the union of finitely many matching sets. ∎
Lemma 17.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$ and $\Phi$ a constraint. Then $\Phi$ is minimal, if, and only if, $\Lambda_{\Phi}$ is compact. More precisely, $\Phi$ is minimal if, and only if, there is a finite matching set $S^{\star}$ such that $\Lambda_{\Phi|S^{\star}}$ is compact. If such $S^{\star}$ exists, $\Lambda_{\Phi|S}$ is compact for any $S\subseteq\mathcal{X}$ that contains $S^{\star}$, and any closed set $S\subseteq\mathcal{X}$ that contains $S^{\star}$ is a matching set.
Proof.
First, notice that if $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ is not minimal, then by Lemma 6 its components are not linearly independent. In particular, there is $\lambda\in\mathbb{R}^{m}$ such that $\lambda\cdot\Phi$ is identicalluy null but $\lambda\neq 0$. Since, for any $\alpha\in\mathbb{R}$, we have that $\alpha\lambda\in\Lambda_{\Phi}$, we conclude that $\Lambda_{\Phi}$ is unbounded. In particular, for any $S^{\star}\subseteq\mathcal{X}$ we also have that $\Lambda_{\Phi|S^{\star}}\supseteq\Lambda_{\Phi}$ is unbounded.
Now, assume that $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ is minimal. Again by Lemma 6, all its components $\Phi_{i}:\mathcal{X}\to\mathbb{R}$ are linearly independent. In particular, by Lemma 20 we can find a set $s$ with finitely many elements, such that $\{\Phi_{1}\big{|}_{s},\dots,\Phi_{m}\big{|}_{s}\}$ is a family of $m$ linearly independent finite-dimensional vectors. By Lemma 14, there is a matching set $S^{\star}\subseteq\mathcal{X}$ that contains $s$ and has finitely many elements. Clearly, we still have that $\{\Phi_{1}\big{|}_{S^{\star}},\dots,\Phi_{m}\big{|}_{S^{\star}}\}$ is a linearly independent family. As $\mathcal{H}_{S^{\star}}$ is a properly constrained hypothesis by Lemma 13, and $\Phi\big{|}_{S^{\star}}$ a constraint for it (Lemma 12), $1-\lambda\cdot\Phi\big{|}_{S^{\star}}$ is an e-variable for $\mathcal{H}_{S^{\star}}$ for every $\lambda\in\Lambda_{\Phi|{S^{\star}}}$, as it is non-negative and has expectation $1$ under every $P\in\mathcal{H}_{S^{\star}}$. In particular, since ${S^{\star}}$ is finite and $\mathcal{H}_{S^{\star}}$ is properly constrained, by Lemma 8 there is a finite constant $a\geq 1$ such that, for any $x\in S^{\star}$ and any $\lambda\in\Lambda_{\Phi|{S^{\star}}}$, we have $1-a\leq\lambda\cdot\Phi(x)\leq 1$. By Lemma 21, $\Lambda_{\Phi|{S^{\star}}}$ is bounded, and so compact by Lemma 16.
Now, let $S\subseteq\mathcal{X}$ be any set that contains $S^{\star}$. Then, $\Lambda_{\Phi|S}$ is compact, as it is a closed (Lemma 16) subset of $\Lambda_{\Phi|{S^{\star}}}$. In particular, $\Lambda_{\Phi}$ is compact. Moreover, from what we have already shown, $\Phi\big{|}_{S^{\star}}$ is a minimal constraint for $\mathcal{H}_{S^{\star}}$, since $\Lambda_{\Phi|S^{\star}}$ is compact. We thus have that $0\in\operatorname*{int}(\operatorname*{conv}\Phi(S^{\star}))\subseteq\operatorname*{int}(\operatorname*{conv}\Phi(S))$. In particular, by Lemma 12, $S$ is compatible, and $\mathcal{H}_{S}$ is a finitely constrained hypothesis on $S$, with $\Phi\big{|}_{S}$ a constraint. Since $\Phi\big{|}_{S}$ contains 00 in the interior of the convex hull of its image, it is a minimal constraint, and so $\mathcal{H}_{S}$ is properly constrained on $S$. By Lemma 13 we conclude that $S$ is a matching set. ∎
Lemma 19.
Let $\mathcal{H}$ be a finitely constrained hypothesis on $\mathcal{X}$. Let $\mathcal{X}_{0}=\bigcup_{P\in\mathcal{H}}\operatorname*{Supp}P$. Then, $\mathcal{X}_{0}$ is closed and $\mathcal{H}$ is a properly constrained hypothesis on $\mathcal{X}_{0}$.
Proof.
Let $\bar{\mathcal{X}}_{0}$ be the closure of $\mathcal{X}_{0}$. Since for every $P\in\mathcal{H}$ we have that $\operatorname*{Supp}P\subseteq\bar{\mathcal{X}}_{0}$, and $\bar{\mathcal{X}}_{0}$ is closed, $\mathcal{H}\subseteq\mathcal{P}_{\bar{\mathcal{X}}_{0}}$ and so it is a hypothesis on $\bar{\mathcal{X}}_{0}$. In particular, $\bar{\mathcal{X}}_{0}$ is a compatible set, and $\mathcal{H}_{\bar{\mathcal{X}}_{0}}=\mathcal{H}\cap\mathcal{P}_{\bar{\mathcal{X}}_{0}}=\mathcal{H}$. Hence, $\mathcal{H}$ is a finitely constrained hypothesis on $\bar{\mathcal{X}}_{0}$ by Lemma 12. By Lemma 9, to show that $\mathcal{H}$ is a properly constrained hypothesis on $\bar{\mathcal{X}}_{0}$ it is enough to show that there is $P^{\star}\in\mathcal{H}$ such that $\operatorname*{Supp}P^{\star}=\bar{\mathcal{X}}_{0}$. As $\operatorname*{Supp}P^{\star}\subseteq\mathcal{X}_{0}\subseteq\bar{\mathcal{X}}_{0}$, this will also yield $\mathcal{X}_{0}=\bar{\mathcal{X}}_{0}$.
Fix a sequence $(x_{i})_{i\geq 1}\subseteq\mathcal{X}_{0}$, dense in $\bar{\mathcal{X}}_{0}$. By construction, for each $i\geq 1$, there is $P_{i}\in\mathcal{H}$ such that $x_{i}\in\operatorname*{Supp}P_{i}$. Let $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ be a constraint for $\mathcal{H}$. Since $P_{i}\in\mathcal{H}$, we have $P_{i}\in\mathcal{P}_{\Phi}$, and so $\zeta_{i}=\langle P_{i},\|\Phi\|_{1}\rangle$ is finite. Let $\alpha_{i}=2^{-i}/(1+\zeta_{i})$. Then, we have $\xi=\sum_{i=1}^{\infty}\alpha_{i}\in(0,1]$. Let $P^{\star}=\sum_{i=1}^{\infty}\frac{\alpha_{i}}{\xi}P_{i}$. Clearly, $P^{\star}\in\mathcal{P}_{\mathcal{X}}$. Moreover, $\langle P^{\star},\|\Phi\|_{1}\rangle=\xi\sum_{i=1}^{\infty}\alpha_{i}\zeta_{i}\leq\xi\sum_{i=1}^{\infty}2^{-i}=\xi$. It follows that $P^{\star}\in\mathcal{P}_{\Phi}$. As $\langle P^{\star},\Phi\rangle=0$, $P^{\star}\in\mathcal{H}$. Moreover, $\operatorname*{Supp}P^{\star}\supseteq\bigcup_{i=1}^{\infty}\operatorname*{Supp}P_{i}\supseteq(x_{i})_{i\geq 1}$. Since $(x_{i})_{i\geq 1}$ is dense in $\bar{\mathcal{X}}_{0}$ and $P^{\star}\in\mathcal{P}_{\bar{\mathcal{X}}_{0}}$ by construction, we deduce that $\operatorname*{Supp}P^{\star}=\bar{\mathcal{X}}_{0}$. ∎
Proposition 5.
Let $\mathcal{H}$ be a properly loosely constrained hypothesis on $\mathcal{X}$, with $\Phi^{\prime}:\mathcal{X}\to\mathbb{R}^{m^{\prime}}$ and $\Phi^{\prime\prime}:\mathcal{X}\to\mathbb{R}^{m^{\prime\prime}}$ tight and slack constraints. Define
[…]
Then, the optimal e-class for $\mathcal{H}$ exists and coincides with
~=$\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}=\big{\{}1-\lambda^{\prime}\cdot\Phi^{\prime}-\lambda^{\prime\prime}\cdot\Phi^{\prime\prime}\;:\;(\lambda^{\prime},\lambda^{\prime\prime})\in\tilde{\Lambda}_{\Phi^{\prime},\Phi^{\prime\prime}}\big{\}}\,.$
Proof.
For $\lambda^{\prime}\in\mathbb{R}^{m^{\prime}}$ and $\lambda^{\prime\prime}\in\mathbb{R}^{m^{\prime\prime}}$, let $E_{\lambda^{\prime},\lambda^{\prime\prime}}=1-\lambda^{\prime}\cdot\Phi^{\prime}-\lambda^{\prime\prime}\cdot\Phi^{\prime\prime}$. It is clear that $\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$ is an e-class. Indeed, for any $(\lambda^{\prime},\lambda^{\prime\prime})\in\tilde{\Lambda}_{\Phi^{\prime},\Phi^{\prime\prime}}$, $E_{\lambda^{\prime},\lambda^{\prime\prime}}\succeq 0$ and, for any $P\in\mathcal{H}$, we have that $\langle P,E_{\lambda^{\prime},\lambda^{\prime\prime}}\rangle=1-\lambda^{\prime}\cdot\langle P,\Phi^{\prime}\rangle-\lambda^{\prime\prime}\cdot\langle P,\Phi^{\prime\prime}\rangle\leq 1$, since all the components of $\lambda^{\prime\prime}$ are non-negative.
Let $\Phi:\mathcal{X}\to\mathbb{R}^{m^{\prime}+m^{\prime\prime}}$ be the continuous function whose first $m^{\prime}$ components are $\Phi^{\prime}$ and the last $m^{\prime\prime}$ are $\Phi^{\prime\prime}$. Since $\mathcal{H}$ is loosely properly constrained, $0\in\operatorname*{rel\,int}(\operatorname*{conv}\Phi)$. Moreover, note that $\mathcal{P}_{\Phi}=\mathcal{P}_{\Phi^{\prime}}\cap\mathcal{P}_{\Phi^{\prime\prime}}$, so =$\mathcal{H}^{\prime}=\{P\in\mathcal{P}_{\Phi^{\prime}}\cap\mathcal{P}_{\Phi^{\prime\prime}}\,:\,\langle P,\Phi\rangle=0\}$ is a properly constrained hypothesis, whereof $\Phi$ is a constraint. By construction $\mathcal{H}^{\prime}\subseteq\mathcal{H}$, so $\mathcal{E}_{\mathcal{H}^{\prime}}\supseteq\mathcal{E}_{\mathcal{H}}$. Moreover, $\mathcal{E}_{\mathcal{H}^{\prime}}^{\vee}\supseteq\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$. In particular, since every $E\in\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$ is a maximal e-variable for $\mathcal{H}^{\prime}$, it must also be a maximal e-variable for $\mathcal{H}$. Hence, by Lemma 2 it is enough to show that $\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$ is a majorising e-class. This can be done reproducing step by step the proofs of Proposition 4 and Theorem 1. Indeed, the preliminary results that were used for those proofs are essentially based on the compactness of $\Lambda_{\Phi|S}$ (for suitable choices of $S\subseteq\mathcal{X}$). This property holds also for $\tilde{\Lambda}_{\Phi^{\prime},\Phi^{\prime\prime}|S}$ (with obvious meaning of notation), as $\tilde{\Lambda}_{\Phi^{\prime},\Phi^{\prime\prime}|S}$ is a closed subset of $\Lambda_{\Phi|S}$ (for the same argument as in Lemma 16). ∎
Appendix B Technical results
B.1 Some linear algebraic results
Lemma 20.
Let $F=\{f_{1},\dots,f_{m}\}$ be a family of $m\geq 1$ functions $\mathcal{X}\to\mathbb{R}$. $F$ is a linearly independent family if, and only if, there is a set $S=\{x_{1},\dots,x_{m}\}\subseteq\mathcal{X}$ such that $\{f_{1}\big{|}_{S},\dots,f_{m}\big{|}_{S}\}$ is a linearly independent family of functions $S\to\mathbb{R}$.
Proof.
If such $S$ exists, then $F$ is an independent family. Indeed, if for some real coefficients $\sum_{i=1}^{m}\alpha_{i}f_{i}=0$, then $\sum_{i=1}^{n}\alpha_{i}f_{i}\big{|}_{S}=0$, which implies that all the coefficients must be null.
We will prove the reverse implication by induction. Obviously, if $m=1$ the statement is true, as $f_{1}$ must be non-zero somewhere for $F$ to be an independent family, and it is enough to set $S=\{x_{1}\}$, with $f(x_{1})\neq 0$. Now, fix $m\geq 2$ and assume that the statement holds for $m-1$. Then, we can find a set $S^{\prime}=\{x_{1},\dots,x_{m-1}\}$ such that the restriction of $\{f_{1},\dots,f_{m-1}\}$ to $S^{\prime}$ is a family that spans the $m-1$ dimensional space of the functions from $S^{\prime}$ to $\mathbb{R}$. In particular, there is a unique vector of coefficients $(\lambda_{1},\dots,\lambda_{m-1})$ such that $f_{m}\big{|}_{S^{\prime}}=\sum_{i=1}^{m-1}\lambda_{i}f_{i}\big{|}_{S^{\prime}}$. Since $f_{m}$ is independent of $\{f_{1},\dots,f_{m-1}\}$, it cannot be that $f_{m}=\sum_{i=1}^{m-1}f_{i}$ on the whole $\mathcal{X}$, and hence there must exist $x_{m}\in\mathcal{X}\setminus S^{\prime}$ such that $f_{m}(x_{m})\neq\sum_{i=1}^{m-1}\lambda_{i}f_{i}(x_{m})$. Let $S=\{x_{1},\dots,x_{m}\}$. Now, let $(\alpha_{1},\dots,\alpha_{m})\in\mathbb{R}^{m}$ be such that $\sum_{i=1}^{m}\alpha_{i}f_{i}\big{|}_{S}=0$. In particular, we have that $\sum_{i=1}^{m-1}(\alpha_{i}+\alpha_{m}\lambda_{i})f_{i}\big{|}_{S^{\prime}}=0$, which implies that for every $i$ from $1$ to $m-1$ it must be that $\alpha_{i}=-\alpha_{m}\lambda_{i}$. Hence, we get that
$0=\sum_{i=1}^{m}\alpha_{i}f_{i}(x_{m})=\alpha_{m}f_{m}(x_{m})-\alpha_{m}\sum_{i=1}^{m-1}\lambda_{i}f_{i}(x_{m})=\alpha_{m}\left(f_{m}(x_{m})-\sum_{i=1}^{m-1}\lambda_{i}f_{i}(x_{m})\right)\,.$
But, we have chosen $x_{m}$ such that $f_{m}(x_{m})-\sum_{i=1}^{m-1}\lambda_{i}f_{i}(x_{m})\neq 0$, and so $\alpha_{m}=0$. It follows that all the coefficients $\alpha_{i}$ are null, and so $\{f_{1}\big{|}_{S},\dots,f_{m}\big{|}_{S}\}$ is an independent family. ∎
Lemma 21.
Let $\{v_{1},\dots,v_{r}\}\subseteq\mathbb{R}^{m}$ be a family of vectors that spans $\mathbb{R}^{m}$. Let $\{(A_{i},B_{i})\}_{i=1}^{r}$ be a family of non-empty bounded intervals. Then, the set
$\Lambda=\{\lambda\in\mathbb{R}^{m}\;:\;\lambda\cdot v_{i}\in[A_{i},B_{i}]\,,\;\forall i=1\dots r\}$
is bounded.
Proof.
Since $\{v_{1},\dots,v_{r}\}$ spans $\mathbb{R}^{m}$, each element $e_{i}$ of the standard basis of $\mathbb{R}^{m}$ can be written as a linear combination $e_{i}=\sum_{j=1}^{r}\alpha_{ij}v_{j}$. Then, if $\lambda\cdot v_{j}$ is in a bounded interval for each $j$, it follows that $\lambda\cdot e_{i}$ is also in a bounded interval. Hence, all the components of $\lambda$ must be bounded if $\lambda\in\Lambda$, and so $\Lambda$ is bounded. ∎
B.2 Some results from convex analysis
First, we state some standard definitions for elementary notions in convex analysis. We refer to Hiriart-Urrut and Lemaréchal (2004) for a thorough introduction to the subject. Given a set $\mathcal{Y}\subseteq\mathbb{R}^{n}$, its convex hull $\mathcal{C}=\operatorname*{conv}\mathcal{Y}$ is defined as the smallest convex set containing $\mathcal{Y}$, or equivalently, as the intersection of all the convex sets that contain $\mathcal{Y}$. The affine hull $\mathcal{A}=\operatorname*{aff}\mathcal{Y}$ of $\mathcal{Y}$ is the smallest affine space in $\mathbb{R}^{n}$ that contains $\mathcal{Y}$, while the linear hull (or span) $\mathcal{L}=\operatorname*{lin}\mathcal{Y}$ of $\mathcal{Y}$ is the smallest vector subspace of $\mathbb{R}^{n}$ containing $\mathcal{Y}$. Clearly,
$\mathcal{Y}\subseteq\mathcal{C}\subseteq\mathcal{A}\subseteq\mathcal{L}\,.$
Let $C\subseteq\mathbb{R}^{n}$ be a convex set and $\mathcal{A}$ its affine hull. The relative interior of $C$ (denoted as $\operatorname*{rel\,int}C$) is the interior of $C$ with respect to the induced topology on $\mathcal{A}$. More explicitly, $x\in C$ is in $\operatorname*{rel\,int}C$ if, and only if, there is an open set $U$ in $\mathbb{R}^{n}$ containing $x$ and such that $U\cap\mathcal{A}\subseteq C$. The relative boundary of $C$ is the set $C\setminus\operatorname*{rel\,int}C$.
Lemma 22 (Carathéodory Theorem).
Let $\mathcal{Y}$ be a set in $\mathbb{R}^{n}$ and $\mathcal{C}$ its convex hull. Denote as $m$ the affine dimension of $\mathcal{Y}$ (namely, the dimension of its affine hull). Let $y\in\mathcal{C}$. Then, there is a set $S\subseteq\mathcal{Y}$ with $m+1$ elements, such that $y\in\operatorname*{conv}S$. Equivalently, there is a Borel probability measure $P\in\mathcal{P}_{\mathcal{Y}}$ such that $\langle P,Y\rangle=y$ and whose support has at most $m+1$ elements.
Proof.
This is a classical elementary result in convex analysis, attributed to Carathéodory. ∎
Corollary 3.
Let $\mathcal{Y}$ be a set in $\mathbb{R}^{n}$ and $\mathcal{C}$ its convex hull. Fix any $y_{0}\in\operatorname*{rel\,int}\mathcal{C}$. For any $y\in\mathcal{Y}$ there is $P\in\mathcal{P}_{\mathcal{Y}}$ whose support has finite cardinality and contains $y$, and such that $\langle P,Y\rangle=y_{0}$.
Proof.
[…]
Lemma 23.
Let $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ be any function. If $0\in\operatorname*{conv}\Phi(\mathcal{X})$, $\dim(\operatorname*{aff}\Phi(\mathcal{X}))=\dim(\operatorname*{Span}\Phi)$.
Proof.
First, note that since $0\in\operatorname*{conv}\Phi(\mathcal{X})$, we have that $0\in\operatorname*{aff}\Phi(\mathcal{X})$, which implies that $\operatorname*{lin}\Phi(\mathcal{X})=\operatorname*{aff}\Phi(\mathcal{X})$. So, it is enough to prove that $\dim(\operatorname*{lin}\Phi(\mathcal{X}))=\dim(\operatorname*{Span}\Phi)$.
Let $m^{\prime}$ be the dimension of $\operatorname*{Span}\Phi$. We can find $m^{\prime}$ components of $\Phi$ that are linearly independent. Without loss of generality, we assume that these are the first $m^{\prime}$ components, and let $\Phi^{\prime}=(\Phi_{1},\dots,\Phi_{m^{\prime}})$, which is a function $\mathcal{X}\to\mathbb{R}^{m^{\prime}}$ with $\operatorname*{Span}\Phi^{\prime}$ of dimension $m^{\prime}$. By Lemma 20, there is a family $S=\{x_{1},\dots,x_{m^{\prime}}\}$ of $m^{\prime}$ distinct points in $\mathcal{X}$ such that the $m^{\prime}\times m^{\prime}$ Gram matrix with components $\Phi_{i}(x_{j})$ (for $i$ and $j$ ranging from $1$ to $m^{\prime}$) has full rank. In particular, $\operatorname*{lin}\Phi^{\prime}(S)=\mathbb{R}^{m^{\prime}}$. Since $\mathbb{R}^{m^{\prime}}\supseteq\operatorname*{lin}\Phi^{\prime}(\mathcal{X})\supseteq\operatorname*{lin}\Phi^{\prime}(S)$, we conclude that $\dim(\operatorname*{lin}\Phi^{\prime}(\mathcal{X}))=m^{\prime}$. So, $\dim(\operatorname*{lin}\Phi(\mathcal{X}))\geq m^{\prime}$. Now, Lemma 20 also implies that no subset of $\mathcal{X}$ with strictly more than $m^{\prime}$ elements can yield a fully ranked Gram matrix. Therefore, there cannot be more than $m^{\prime}$ independent vectors in $\operatorname*{lin}\Phi(\mathcal{X})$, and so $\dim(\operatorname*{lin}\Phi(\mathcal{X}))=m^{\prime}$. ∎
Lemma 24.
Let $\mathcal{X}$ be a closed set in $\mathbb{R}^{n}$ and $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ a Borel function. Let $\mathcal{C}=\operatorname*{conv}\Phi(\mathcal{X})$, and $\partial\mathcal{C}$ its relative boundary. If there is $P\in\mathcal{P}_{\Phi}$ such that $\langle P,\Phi\rangle=0$, then $0\in\mathcal{C}$. Moreover, if $0\in\partial\mathcal{C}$, then any $P\in\mathcal{P}_{\Phi}$ such that $\langle P,\Phi\rangle=0$ satisfies $\operatorname*{Supp}P\subseteq\Phi^{-1}\big{(}\bigcap_{\pi\in\Pi_{0}}\pi\big{)}$, where $\Pi_{0}$ is the set of all supporting hyperplanes of $\mathcal{C}$ containing 00, and so in particular $\operatorname*{Supp}P\subseteq\Phi^{-1}(\partial\mathcal{C})$.
Proof.
We will show the first statement by induction on $m$. If $m=0$, $\mathcal{C}=\{0\}=\mathbb{R}^{0}$, and the result is trivial. Now, for $m\geq 1$, assume that the result holds for $m-1$. We prove by contradiction that it must hold for $m$. Assume that $0\notin\mathcal{C}$. By the hyperplane separating theorem for convex sets, there is a vector $v\in\mathbb{R}^{m}$ such that $v\cdot y\geq 0$, for every $y\in\mathcal{C}$. Since $\Phi(\mathcal{X})\subseteq\mathcal{C}$, any $P$ such that $\langle P,\Phi\rangle=0$ must be supported on the set $\mathcal{X}_{0}=\{x\in\mathcal{X}\,:\,\Phi(x)\cdot v=0\}$. Since $\mathcal{X}_{0}\cong\mathbb{R}^{m-1}$, we conclude that $0\in\operatorname*{conv}\Phi(\mathcal{X}_{0})$ by the inductive hypothesis, which is a contradiction since $\operatorname*{conv}\Phi(\mathcal{X}_{0})\subseteq\mathcal{C}$.
For the second claim, assume that $0\in\partial\mathcal{C}$ and fix any $P\in\mathcal{P}_{\Phi}$ such that $\langle P,\Phi\rangle=0$. $\Pi_{0}$ is non-empty as $0\in\partial\mathcal{C}$, and for any supporting hyperplane $\pi\in\Pi_{0}$ there is a linear map $\phi$ such that $\pi=\{\phi=0\}$ and $\Phi(\mathcal{X})\subseteq\mathcal{C}\subseteq\{\phi\geq 0\}$. Hence, $\operatorname*{Supp}P\subseteq\Phi^{-1}(\pi)=\Phi^{-1}(\pi\cap\Phi(\mathcal{X}))\subseteq\Phi^{-1}(\partial\mathcal{C})$, since $\pi\cap\Phi(\mathcal{X})\subseteq\partial\mathcal{C}$. As this is true for every $\pi\in\Pi_{0}$ we also have that $\operatorname*{Supp}P\subseteq\Phi^{-1}\big{(}\bigcap_{\pi\in\Pi_{0}}\pi\big{)}$. ∎
Lemma 25.
Let $\mathcal{X}$ be a set with finite cardinality $d\geq 1$. Then, the relative interior of $\mathcal{P}_{\mathcal{X}}$ (which can be seen as a subset of $\mathbb{R}^{d}$) is the set $\{P\in\mathcal{P}_{\mathcal{X}}\,:\,\operatorname*{Supp}P=\mathcal{X}\}$.
Proof.
Let $\mathbb{R}^{d}_{\geq 0}$ denote the subset of $\mathbb{R}^{d}$ of vectors whose components are all non-negative. Its relative interior is >$\mathbb{R}^{d}_{>0}$, the set of vectors with only strictly positive components. Let $\mathbf{1}\in\mathbb{R}^{d}$ denote the vector with all components equal to one and $V=\{v\in\mathbb{R}^{d}\,:\,v\cdot\mathbf{1}=1\}$, whose relative interior is $V$ itself. Then $\mathcal{P}_{\mathcal{X}}=\mathbb{R}^{d}_{\geq 0}\cap V$, which is the non-empty intersection of two convex sets. Since $\operatorname*{rel\,int}V\cap\operatorname*{rel\,int}\mathbb{R}^{d}_{\geq 0}\neq\varnothing$, from Proposition 2.1.10 in Chapter A of Hiriart-Urrut and Lemaréchal, 2004 we get
$\operatorname*{rel\,int}\mathcal{P}_{\mathcal{X}}=\operatorname*{rel\,int}V\cap\operatorname*{rel\,int}\mathbb{R}^{d}_{\geq 0}=V\cap\mathbb{R}^{d}_{>0}=\mathcal{P}_{\mathcal{X}}\cap\mathbb{R}^{d}_{>0}\,,$
which is precisely the set of fully supported probability measures on $\mathcal{X}$. ∎
Lemma 26.
Let $V\subseteq\mathbb{R}^{d}$ be a linear subspace. Denote as $\Delta$ the simplex in $\mathbb{R}^{d}$ and as $\Delta^{\circ}$ its relative interior. If $V\cap\Delta^{\circ}\neq\varnothing$, we have that $\operatorname*{lin}(V\cap\Delta)=V$.
Proof.
Everything is trivial if $d=0$, so assume $d\geq 1$. Let $v\in V\cap\Delta^{\circ}\neq\varnothing$. As $v\in\Delta^{\circ}$, by Lemma 25 all its component are strictly positive, so there is an open ball $B\subset\mathbb{R}^{d}_{\geq 0}$ centred at $v$. It follows that
$V=\operatorname*{lin}V\supseteq\operatorname*{lin}(V\cap\Delta)=\operatorname*{lin}(V\cap\mathbb{R}^{d}_{\geq 0})\supseteq\operatorname*{lin}(V\cap B)\,.$
However, since $v\in V$, $\operatorname*{lin}(V\cap B)=\operatorname*{lin}(V\cap B_{0})=V$, where $B_{0}=B-v$ is the open ball with same radius as $B$ and centred in 00. So, we conclude. ∎
Appendix C An example of non-existence of the optimal e-class
Let $\mathcal{X}=[0,1]$. Denote as $V$ any non-Borel set in $[0,1]$, which contains 00 and does not contain $1$. For $x\in[0,1]$, let $P_{x}$ denote $\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{x}$. Define $\mathcal{H}$ as follows
=$\mathcal{H}=\big{\{}P_{x}\;:\;x\in[0,1]\big{\}}\cup\mathcal{P}_{V}\,.$
Recall that $\mathcal{P}_{V}$ denote the set of all probability measures in $[0,1]$ whose support lies in $V$. So, even if $V$ is non-measurable, $\mathcal{P}_{V}$ is well defined. Clearly, $\mathcal{H}$ is a hypothesis on $[0,1]$, and
$\sup_{P\in\mathcal{H}}P(\{x\})=\begin{cases}1&\text{if $x\in V$;}\\ 1/2&\text{otherwise.}\end{cases}$
Now, define the function $A:[0,1]\to\mathbb{R}$ as $A(x)=1/\sup_{P\in\mathcal{H}}P(\{x\})$. Note that $A$ is not Borel measurable, so $A\notin\mathcal{E}_{\mathcal{H}}$. For any $E\in\mathcal{E}_{\mathcal{H}}$ and $x\in[0,1]$, we have
$E(x)\leq A(x)E(x)\sup_{P\in\mathcal{H}}P(\{x\})\leq A(x)\sup_{P\in\mathcal{H}}\langle P,E\rangle\leq A(x)\,.$
We will show that every e-variable $E$ such that $E(1)=2$ cannot be maximal. Since every majoriser $\hat{E}$ of such $E$ has to still satisfy $\hat{E}(1)=2$ (as $22=E(1)\leq\hat{E}(1)\leq A(1)=$ 2), we will then conclude that $E$ does not have any maximal majoriser and so that the class of all maximal e-variables is not a majorising class. Lemma 2 will yield that the optimal e-class does not exist.
So, fix an e-variable $E$ such that $E(1)=2$ (such an e-variable exists, for instance one can pick the function that is $2$ on $1$ and 00 everywhere else). Since $\langle P_{1},E\rangle\leq 1$, it must be that $E(0)=0$. The restriction of $A$ to $(0,1]$ is non-Borel, so there must be $x_{0}\in(0,1]$ such that $E(x_{0})<A(x_{0})$. Define $E^{\prime}:[0,1]\to[0,+\infty)$ as $E^{\prime}(x)=E(x)$, for $x\neq x_{0}$, and $E^{\prime}(x_{0})=A(x_{0})$. Since $E^{\prime}\preceq A$, we have that $E^{\prime}(x)\leq 1$ for any $x\in V$. In particular, for any $P\in\mathcal{P}_{V}$ we have $\langle P,E^{\prime}\rangle\leq 1$. Moreover, $E^{\prime}(0)=E(0)=0$ and $E^{\prime}(x)\preceq 2$ for all $x\in[0,1]$. Thus, $\langle P_{x},E^{\prime}\rangle=\frac{1}{2}E^{\prime}(x)\leq 1$, for any $x\in[0,1]$. We deduce that $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$. Since $E^{\prime}\succ E$, $E$ is not maximal, and so we conclude.
Appendix D Proof of Proposition 1
Lemma 27.
In $\mathbb{R}^{m}$, a straight line $r$ intersecting the interior of a convex set $C$ is fully contained in $C$ if, and only if, it is parallel to all the supporting hyperplanes of the closure of $C$.
Proof.
First, note that since $r$ intersects the interior of $C$, it intersects the supporting hyperplanes of the closure of $C$ in at most one point, and hence cannot lie on one of these hyperplanes. Now, if it intersects one of these hyperplanes in only one point, then it cannot be fully contained in $C$. On the other hand, if $r$ is not contained in $C$, $r$ must intersect the boundary of $C$. Since every point of the boundary of $C$ is contained in at least one supporting hyperplane of the closure of $C$, we conclude that $r$ must intersect this hyperplane. ∎
Lemma 28.
Let $C\subseteq\mathbb{R}^{m}$ be a convex set with $0\in\operatorname*{int}C$. Let $V=\{v\in\mathbb{R}^{m}\,:\,\operatorname*{lin}\{v\}\subseteq C\}$. Then, $V$ is a subspace of $\mathbb{R}^{m}$. Moreover, there is a (possibly non orthogonal) basis $\mathcal{B}$ of $V^{\perp}$ such that $\inf_{w\in C}v\cdot w>-\infty$, for all $v\in\mathcal{B}$.
Proof.
Clearly, $0\in V$. Now, let $v$ and $v^{\prime}$ belong to $V$. Since $C$ is convex, the fact that both $\operatorname*{lin}\{v\}$ and $\operatorname*{lin}\{v^{\prime}\}$ lie in $C$ implies that $\operatorname*{lin}\{v,v^{\prime}\}\subseteq C$. In particular, for any real coefficients $\alpha$ and $\alpha^{\prime}$, we have $\alpha v+\alpha^{\prime}v^{\prime}\in V$. Hence, $V$ is a subspace of $\mathbb{R}^{m}$.
Let $S$ denote the set of all the normal vectors to the supporting hyperplanes of the closure of $C$, and let $U=\operatorname*{lin}S$. Since $0\in\operatorname*{int}C$, by Lemma 27, $U=V^{\perp}$. By definition of $U$, it is hence possible to find a basis of $V^{\perp}$ made of elements in $S$. Now, let $v\in S$. Then there is a supporting hyperplane $\pi$ of the closure of $C$ that is orthogonal to $v$. This means that there is $\alpha\in\mathbb{R}$ such that $\pi=\{w\in\mathbb{R}^{m}\,:\,v\cdot w=\alpha\}$. Since the origin is in the interior of $C$, it must be that $\alpha\neq 0$. If $\alpha<0$, then $C\subseteq\{w\in\mathbb{R}^{m}\,:\,v\cdot w\geq\alpha\}$, by definition of supporting hyperplane and since $0\in C$. In such a case it is clear that $\inf_{w\in C}v\cdot w\geq\alpha>-\infty$. On the other hand, if $\alpha>0$, we replace $v$ with $-v$ in the basis, and we go back to the case $\alpha<0$. ∎
Lemma 29.
Let $\mathcal{Y}\subseteq\mathbb{R}^{m}$ be any set. Let $C$ denote the convex hull of $\mathcal{Y}$. Assume that for some $\varepsilon>0$, the closed 1-ball of radius $\varepsilon$ centred in 00 is included in $C$. Define $L:\mathbb{R}^{m}\to(0,1]$ as
$L(y)=\frac{\varepsilon}{\varepsilon+\|y\|_{1}}\,.$
Let $V=\{v\in\mathbb{R}^{m}\,:\,\operatorname*{lin}\{v\}\subseteq C\}$ and $\Pi$ be the orthogonal projection from $\mathbb{R}^{m}$ to $V^{\perp}$. For any $y\in\mathbb{R}^{m}$, let $\mathcal{P}_{y}^{\circ}$ denote the set of measures in $\mathcal{P}_{\mathcal{Y}\cup\{y\}}$ whose support has finitely many elements, and let $\mathcal{H}_{y}^{\circ}=\{P\in\mathcal{P}_{y}^{\circ}\,:\,\langle P,{Y}\rangle=0\}$. Then, $\sup_{P\in\mathcal{H}_{y}^{\circ}}P(\{y\})\geq L(\Pi(y))$, for any $y\in\mathbb{R}^{m}$.
Proof.
For $y\in\mathbb{R}^{m}$, let $S_{y}=\sup_{P\in\mathcal{H}_{y}^{\circ}}P(\{y\})$. We will proceed in two steps.
  1. 1.
    First, we show that $S_{y}\geq L(y)$, for any $y\in\mathbb{R}^{m}$.
  2. 2.
    Then, we prove the tighter bound $S_{y}\geq L(\Pi(y))$, for any $y\in\mathbb{R}^{m}$.
For $y=0$, these claims are trivially true (since $\delta_{0}\in\mathcal{H}_{0}^{\circ}$). We will henceforth assume that $y\neq 0$.
  1. 1.
    Fix $y\in\mathbb{R}^{m}$ non-zero. Since the centred closed ball of radius $\varepsilon$ is in $C$, $z=-\varepsilon y/\|y\|_{1}\in C$. Let $\mathcal{P}_{y}^{\circ}$ denote the set of probability measures on $\mathbb{R}^{m}$ whose support has finitely many elements and is included in $\mathcal{Y}$. By Lemma 22, there is $P\in\mathcal{P}_{y}^{\circ}$, such that $\langle P,Y\rangle=z$. Now consider $P^{\prime}=\frac{\varepsilon}{\varepsilon+\|y\|_{1}}\delta_{y}+\frac{\|y\|_{1}}{\varepsilon+\|y\|_{1}}P$. Clearly, $P^{\prime}\in\mathcal{P}_{y}^{\circ}$. Moreover, it is easily checked that $\langle P^{\prime},Y\rangle=0$, so $P^{\prime}\in\mathcal{H}_{y}^{\circ}$. Since $P^{\prime}(\{y\})\geq L(y)$, it follows that $S_{y}\geq L(y)$.
  2. 2.
    Fix a non-zero $y$ in $\mathbb{R}^{m}$. By Lemma 28, $V$ is a subspace. Let $y_{\perp}=\Pi(y)$ and $y_{V}=y-y_{\perp}$. Then, $y_{V}\in V$. Thus, for any $u>0$ we have that $z_{u}=-uy_{V}\in C$. By Lemma 22, there is $P_{u}\in\mathcal{P}_{y}^{\circ}$ such that $\langle P_{u},Y\rangle=z_{u}$. Let $y_{u}=\frac{u}{1+u}y_{\perp}$ and $P^{\prime}_{u}=\frac{u}{1+u}\delta_{y}+\frac{1}{1+u}P_{u}\in\mathcal{P}_{y}^{\circ}$. Clearly, $\langle P_{u}^{\prime},Y\rangle=\frac{u}{1+u}(y-y_{V})=y_{u}$. Fix an arbitrary $\alpha>0$. From what we have already shown, we know that there is $P^{\prime\prime}_{u}\in\mathcal{H}_{y_{u}}^{\circ}$ such that $P^{\prime\prime}_{u}(\{y_{u}\})>L(y_{u})-\alpha$. Denoting $p=P^{\prime\prime}_{u}(\{y_{u}\})$, let $P^{\prime\prime}_{u}=p\delta_{y_{u}}+(1-p)P^{\prime}$, with $P^{\prime}\in\mathcal{P}_{y}^{\circ}$. Let $\hat{P}_{u}=pP^{\prime}_{u}+(1-p)P^{\prime}$. Then, $\hat{P}_{u}\in\mathcal{P}_{y}^{\circ}$ and $\langle\hat{P}_{u},Y\rangle=py_{u}+(1-p)\langle P^{\prime},Y\rangle=\langle P^{\prime\prime}_{u},Y\rangle=0$, since $P^{\prime\prime}_{u}\in\mathcal{H}_{y_{u}}^{\circ}$. So, $\hat{P}_{\mu}\in\mathcal{H}_{y}^{\circ}$ and $\hat{P}_{u}(\{y\})\geq pP^{\prime}_{u}(\{y\})\geq\frac{up}{1+u}>\frac{u}{1+u}(L(y_{u})-\alpha)$. The whole construction is valid for any $u>0$ and $\alpha>0$. So, $S_{y}\geq\lim_{\alpha\to 0}\lim_{u\to+\infty}\frac{u}{1+u}(L(y_{u})-\alpha)=L(\Pi(y))$, since $L$ is continuous and $y_{u}\to y_{\perp}=\Pi(y)$ for $u\to+\infty$.
This concludes the proof. ∎
Lemma 30.
Let $\mathcal{Y}\subseteq\mathbb{R}^{m}$ be any set. Let $C$ denote the convex hull of $\mathcal{Y}$, and assume that $0\in\operatorname*{int}C$. Let $\mathcal{P}_{1}=\{P\in\mathcal{P}_{\mathcal{Y}}\,:\,\langle P,\|Y\|_{1}\rangle\leq 1\}$, and =$\mathcal{H}=\{P\in\mathcal{P}_{1}\,:\,\langle P,Y\rangle=0\}$. Let $\mathcal{H}^{\circ}$ denote the set of measures in $\mathcal{H}$ whose support has finite cardinality. Then, there is a continuous function $A:\mathcal{Y}\to[1,+\infty)$, such that, for any $y\in\mathcal{Y}$, $A(y)\sup_{P\in\mathcal{H}^{\circ}}P(\{y\})\geq 1$, and $\langle P,A\rangle<+\infty$ for all $P\in\mathcal{H}$.
Proof.
Since $0\in\operatorname*{int}C$, we can fix $\varepsilon>0$ such that the centred closed 1-ball of radius $\varepsilon$ is in $C$. Let $L$, $V$, and $\Pi$ be as in Lemma 29. For $y\in\mathcal{Y}$, let $A(y)=1/L(\Pi(y))=1+\|\Pi(y)\|_{1}/\varepsilon$. Then, $A:\mathcal{Y}\to[1,+\infty)$ is continuous and, for any $y\in\mathcal{Y}$, $A(y)\sup_{P\in\mathcal{H}^{\circ}}P(\{y\})\geq 1$ by Lemma 29.
Let $p\leq m$ be the dimension of $V^{\perp}$. By Lemma 28, we can fix a basis =$\mathcal{B}=\{v_{1},\dots,v_{m}\}$ of $\mathbb{R}^{m}$, such that $\{v_{p+1},\dots v_{m}\}$ is a basis of $V$ and $\{v_{1},\dots,v_{p}\}$ is a basis of $V^{\perp}$, where for each $i\leq p$ we have that $\inf_{w\in C}w\cdot v_{i}>-\infty$. For $y\in\mathbb{R}^{m}$, let $\|y\|_{\mathcal{B}}=\sum_{i=1}^{m}|v_{i}\cdot y|$. Since $\mathcal{B}$ is a basis, this is a norm on $\mathbb{R}^{m}$. As all norms on finite dimensional spaces are equivalent, we can find a constant $K>0$ such that, for any $y\in\mathbb{R}^{m}$, $\|y\|_{1}\leq K\|y\|_{\mathcal{B}}$. It follows that, for any $y\in\mathcal{Y}$,
$A(y)\leq 1+\frac{K}{\varepsilon}\|\Pi(y)\|_{\mathcal{B}}=1+\frac{K}{\varepsilon}\sum_{i=1}^{p}|v_{i}\cdot y|\,.$
Now, for any $i\leq p$ and any $P\in\mathcal{H}$, we have that $\langle P,v_{i}\cdot Y\rangle=0$. Since $v_{i}\cdot Y$ is bounded from below (as $Y$ is supported in $C$), it follows that $\langle P,|v_{i}\cdot Y|\rangle$ is finite. Hence, $\langle P,A\rangle<+\infty$. ∎
Proposition 1.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$. There is a continuous function $A:\mathcal{X}\to[1,+\infty)$ such that $\langle P,A\rangle<+\infty$ for every $P\in\mathcal{H}$, $\sup_{P\in\mathcal{H}}P(\{x\})\geq 1/A(x)$ for all $x\in\mathcal{X}$, and $E\preceq A$ for any $E\in\mathcal{E}_{\mathcal{H}}$.
Proof.
By Lemma 5, we can find a minimal constraint $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ for $\mathcal{H}$. Let $\mathcal{Y}=\Phi(\mathcal{X})$ and $C=\operatorname*{conv}\mathcal{Y}$. Define =$\mathcal{H}^{\prime}=\{\Phi^{\#}P\,:\,P\in\mathcal{H}\}$, where $\Phi^{\#}P$ is the push-forward of $P$ under $\Phi$. Let $\mathcal{H}^{\circ}$ be the set of measures in $\mathcal{P}_{\mathcal{Y}}$ with zero mean and finite support, which is non-empty by Lemma 22. Fix any $P^{\circ}\in\mathcal{H}^{\circ}$. For each $y\in\operatorname*{Supp}P^{\circ}$, $\xi_{y}=\Phi^{-1}(\{y\})$ is Borel measurable. Since all these $\xi_{y}$ are disjoint, we can define a Borel measure $\tilde{P}^{\circ}$ on $\mathcal{X}$, supported on $\cup_{y\in\operatorname*{Supp}P^{\circ}}\xi_{y}$, such that $\tilde{P}^{\circ}(\xi_{y})=P^{\circ}(\{y\})$. It’s easy to verify that $P^{\circ}=\Phi^{\#}\tilde{P}^{\circ}$ and $\tilde{P}^{\circ}\in\mathcal{P}_{\Phi}$. So, $\langle\tilde{P}^{\circ},\Phi\rangle=\langle P^{\circ},Y\rangle=0$. Hence, $\tilde{P}^{\circ}\in\mathcal{H}$ and $P^{\circ}\in\mathcal{H}^{\prime}$. We conclude that $\mathcal{H}^{\circ}\subseteq\mathcal{H}^{\prime}$.
Since $\Phi$ is a minimal constraint, $0\in\operatorname*{int}C$. Since for every $P^{\prime}\in\mathcal{H}^{\prime}$ we have that $\langle P^{\prime},Y\rangle=0$ and $\operatorname*{Supp}P^{\prime}\subseteq\mathcal{Y}$, by Lemma 30 there is a continuous $\tilde{A}:\mathcal{Y}\to[1,+\infty)$, such that $\langle P^{\prime},\tilde{A}\rangle<+\infty$ for all $P^{\prime}\in\mathcal{H}^{\prime}$, and $\sup_{P^{\prime}\in\mathcal{H}^{\prime}}P^{\prime}(\{y\})\tilde{A}(y)\geq\sup_{P^{\prime}\in\mathcal{H}^{\circ}}P^{\prime}(\{y\})\tilde{A}(y)\geq 1$, for any $y\in\mathcal{Y}$. Now, let $A=\tilde{A}\circ\Phi$. Clearly, $A$ is a continuous function from $\mathcal{X}$ to $[1,+\infty)$. Moreover, for any $P\in\mathcal{H}$ we have $\langle P,A\rangle=\langle P,\tilde{A}\circ\Phi\rangle=\langle\Phi^{\#}P,\tilde{A}\rangle<+\infty$, since $\Phi^{\#}P\in\mathcal{H}^{\prime}$.
Now, fix $x\in\mathcal{X}$. Let $y=\Phi(x)$ and $\xi=\Phi^{-1}(\{y\})\subseteq\mathcal{X}$. Clearly, $\xi$ is Borel measurable, as $\Phi$ is continuous. Fix any $P^{\prime}\in\mathcal{H}^{\prime}$. By construction, there is $P\in\mathcal{H}$ such that $P^{\prime}=\Phi^{\#}P$. In particular, $P(\xi)=P^{\prime}(\{y\})$. Now, let $P^{\star}$ be the non-negative measure on $\mathcal{X}$ that is equal to $P$ on $\mathcal{X}\setminus\xi$, and gives no mass to $\xi$. Because $P\in\mathcal{H}$, $\int_{\mathcal{X}}\|\Phi\|_{1}\mathrm{d}P^{\star}<+\infty$ and $\int_{\mathcal{X}}\Phi\mathrm{d}P^{\star}=-P(\xi)y$. Let $\hat{P}=P^{\star}+P(\xi)\delta_{x}$. Then, $\hat{P}\in\mathcal{P}_{\Phi}$. Moreover, $\langle\hat{P},\Phi\rangle=P(\xi)y+\langle P^{\star},\Phi\rangle=0$. Hence, $\hat{P}\in\mathcal{H}$, and $\hat{P}(\{x\})=P(\xi)=P^{\prime}(\{y\})$. Thus, for any $P^{\prime}\in\mathcal{H}^{\prime}$, there is $\hat{P}\in\mathcal{H}$ such that $\hat{P}(\{x\})=P^{\prime}(\{y\})$. So, $\sup_{P\in\mathcal{H}}P(\{x\})\geq\sup_{P^{\prime}\in\mathcal{H}^{\prime}}P^{\prime}(\{\Phi(x)\})\geq 1/\tilde{A}(\Phi(x))=1/A(x)$.
Finally, for any e-variable $E$ and any $x\in\mathcal{X}$, we have
$E(x)\leq A(x)\sup_{P\in\mathcal{H}}P(\{x\})E(x)\leq A(x)\sup_{P\in\mathcal{H}}\langle P,E\rangle\leq A(x)\,,$
and so we conclude. ∎